### v4

# Nellie: A Neuro-Symbolic Inference Engine for Grounded, Compositional, and Explainable Reasoning

Abstract
Our goal is a modern approach to answering questions via systematic reasoning where answers are supported by human interpretable proof trees grounded in an NL corpus of authoritative facts. Such a system would help alleviate the challenges of interpretability and hallucination with modern LMs, and the lack of grounding of current explanation methods (e.g., Chain-of-Thought). This paper proposes a new take on Prolog-based inference engines, where we replace handcrafted rules with a combination of neural language modeling, guided generation, and semiparametric dense retrieval. Our implementation, Nellie, is the first system to demonstrate fully interpretable, end-to-end grounded QA as entailment tree proof search, going beyond earlier work explaining known-to-be-true facts from text. In experiments, Nellie outperforms a similar-sized state-of-the-art reasoner (Tafjord et al., 2022) while producing knowledge-grounded explanations. We also find Nellie can exploit both semi-structured and NL text corpora to guide reasoning. Together these suggest a new way to jointly reap the benefits of both modern neural methods and traditional symbolic reasoning.
1 Introduction
Large language models (LLMs) are impressive at question-answering (QA), but it remains challenging to understand how answers systematically follow from authoritative information. This general opacity is a growing impediment to widespread use of LLMs, e.g., in critical applications such as medicine, law, and hiring decisions, where interpretability and trust are paramount. While there has been rapid progress on having LLMs generate systematic explanations for their answers, e.g., Chain-of-Thought (Wei et al., 2022), Entailer (Tafjord et al., 2022), or Maieutic Prompting (Jung et al., 2022), such explanations are not grounded in external facts and may include hallucinations (Ji et al., 2022). Rather, what would be desirable - and what this work pursues - is a system that systematically reasons over authoritative text: to support answers with human interpretable proof trees grounded in the text, while not requiring translation to an entirely symbolic formalism.
Figure 1: Given a query, Nellie performs a neuro-symbolic backward chaining search for proof trees whose leaves are grounded in a corpus of facts. First, it generates candidate decomposition rules guided by both retrieved facts and templates. Then, it recursively tries to prove the rules’ conditions via entailment from the corpus or further decomposition.
Our approach is to revisit the behavior of expert systems (Jackson, 1986; Metaxiotis et al., 2002). Expert systems are appealing for their explainable behavior: decisions are made by constructing a well-formed symbolic proof from explicit, formally represented knowledge authored by a knowledge engineer in consultation with a domain expert. However, as expert systems are known to be both expensive and brittle (Musen and Van der Lei, 1988), the AI community has turned to neuro-symbolic mechanisms that use large language models (LLMs) to reason over natural language (NL). Reasoning over NL does not require a symbolic formalism, and allows the use of inferentially powerful LLMs, but also loses the systematic, verifiable proofs that expert systems produced. This motivates our pursuit of a new way to jointly reap the benefits of both modern neural methods and traditional reasoning.
We desire the following expert system-inspired desiderata:
  1. 1.
    Grounding inferences fully and scalably in a corpus of knowledge from an authoritative human source.
  2. 2.
    Logical direction, showing how a given hypothesis (and all intermediate inferences leading up to it) follows as the logical consequence of the underlying knowledge source
  3. 3.
    Competitive end-to-end QA performance in a complex domain requiring diverse forms of reasoning.
As illustrated in Figure 2, various eXplainable QA (XQA) methods accomplish 2 of these criteria, but not all 3. Some methods, like Chain-of-Thought, extract inference chains or logically structured graphs fully from a LLM without grounding in verified knowledge; others, like ExplanationLP (Thayaparan et al., 2021), compose graphs of grounded facts, but do not show logical direction; finally, LAMBADA (Kazemi et al., 2023) achieves both direction and grounding, but is limited to simple domains with provided NL Horn rules and small sets of facts (1-2 dozen). In contrast, we desire a system that handles a much larger corpus (10K statements) and doesn’t need handcrafted rules.
Figure 2: Comparison of approaches to neural XQA. Each approach leads to NL graphs in support of a query. Our proposal is to produce logically directed explanations containing model-generated intermediate steps while grounding a tree in verified facts without relying on handwritten horn clauses.
To achieve all three desiderata, we reuse the general inference framework of an expert system, but replace handcrafted rules with a combination of neural language modeling, guided generation, and semiparametric dense retrieval. We demonstrate this in a system called Nellie, the Neuro-Symbolic Large LM Inference Engine. Nellie leverages LLMs as proposal functions in the search of proof trees showing how a conclusion follows via entailment from an external corpus of NL statements. The “symbols” that our neuro-symbolic engine reasons over are free-form NL sentences. Rather than require a knowledge engineer to carefully write hundreds of inference rules as in the classical setting, Nellie employs an LLM as a dynamic rule generator  (DRG; Kalyanpur et al., 2021) to generate (rather than retrieve) candidate rules that decompose a hypothesis into subqueries that, if themselves proved recursively, will prove that hypothesis via composition (Figure 1). In this way, Nellie can answer the question posed by the classical expert system: “Does this statement systematically follow from my knowledge, or not?”.
Nellie is built upon a backward chaining symbolic theorem prover written in Prolog, implemented using a few meta-rules specifying how inference should proceed. These include checking for entailment of a hypothesis against a retrieved fact, or decomposing it into a conjunction of subqueries to recursively prove. To treat NL sentences as if they were symbols in a purely symbolic search algorithm, we use an NLI-based form of weak unification (Sessa, 2002; Weber et al., 2019) between a queried statement and a corpus fact. This leads to interpretable proofs similar to the compositional entailment trees of prior work (e.g., EntailmentBank (Dalvi et al., 2021)), while tackling the additional challenge of multiple-choice QA.
Nellie is designed to search across hundreds of trees whose leaves come from one of two types of knowledge: (a) a corpus of semi-structured text statements, such as NL renditions of database entries, or (b) a corpus of free-form NL sentences. Many correct proofs might exist for a given hypothesis, but much fewer will be fully groundable in the provided corpus, making the search harder than for ungrounded alternatives like Entailer (Tafjord et al., 2022) or COMET-Dynagen (Bosselut et al., 2021). To improve grounding, we introduce two guiding methods to boost the likelihood of generating rules whose conditions match against the available text: (1) For applications where a semi-structured corpus is available, Nellie leverages this structure via templates to help steer rule generation towards syntax that is more likely to match corpus entries. (2) For both free-form and semi-structured corpora, Nellie retrieves and conditions on statements to help steer generation towards trees grounded in them. Ablation experiments show these together and individually improve Nellie’s reasoning.
Nellie expands upon work explaining known-to-be-true facts from text (SCSEARCH; Bostrom et al., 2022),(IRGR; Ribeiro et al., 2022) to perform QA, including searching for and comparing trees for conflicting options (both true and false). Experiments find Nellie outperforms a similar-sized SOTA reasoner (Entailer) while producing compositional trees showing how all decisions (top-level and intermediate) are grounded in corpus facts. Our contributions are thus:
  1. 1.
    An architecture for systematic reasoning from a corpus of textual knowledge to answer questions. This can be viewed as a modern approach to expert system inference, but without requiring a formal knowledge representation.
  2. 2.
    An implementation, Nellie, [^1] Code will be released at https://github.com/JHU-CLSP/NELLIE that outperforms a similar-sized SOTA reasoner (Entailer-3B) while producing grounded trees. To our knowledge, this is the first system to perform grounded xQA as NL entailment tree search.
2 Related Work
Theorem Proving over Language A long-standing approach to reasoning over NL is to project it into a symbolic form, such as for QA (Green et al., 1961; Zelle and Mooney, 1996) or entailment (Bos and Markert, 2005). Provided the translation from NL to symbolic representation is accurate, one can leverage fast and scalable solutions for discrete symbolic inference (Riazanov and Voronkov, 2002; Kautz et al., 1992; Kautz and Selman, 1999). However, reliably translating broad-domain NL into an adequately expressive formalism for reasoning is a challenge (Schubert, 2015).
Recent work explores methods of handling NL without needing to parse it to another formalism. Variants of neural theorem provers (NTPs; Rocktäschel and Riedel, 2017) such as NLProlog (Weber et al., 2019) reconcile NL facts with symbolic reasoning by learning embeddings for both the facts and the symbols in a theorem prover, then using weak unification to backward chain. (Kalyanpur et al., 2021) inject neural reasoning into a Boxer/Prolog-based symbolic reasoner via a special LM-calling predicate. (Arabshahi et al., 2021b) combine handwritten symbolic rules with neural symbol embeddings in order to classify conversational intents. Other work explores using LMs to emulate stepwise (Tafjord et al., 2021; Kazemi et al., 2023) or end-to-end (Clark et al., 2021; Picco et al., 2021) logical reasoning over small rulebases converted to NL.
Systematic Explanation Generation Recent works have used LMs to generate NL reasoning graphs in support of an answer. “Chain-of-Thought” prompting (Wei et al., 2022; Kojima et al., 2022), elicits free-form inference hops from the LM before it generates an answer. Other text graph generation methods connect model-generated statements via common sense relations (Bosselut et al., 2021; Arabshahi et al., 2021a) or for/against influence (Madaan et al., 2021; Jung et al., 2022), though these are not knowledge-grounded and don’t address entailment (see Figure 2).
The EntailmentBank dataset (Dalvi et al., 2021) has driven research towards the construction of explanation trees, challenging models to produce stepwise entailment proofs of a statement using a set of provided support facts. This direction builds upon works on explainable reasoning that build graphs from KB-retrieved sets of support statements, but stop short of showing their role in logical entailment (Pan et al., 2019; Jansen and Ustalov, 2019; Yadav et al., 2019; Valentino et al., 2022). Components of our framework are related to concurrent approaches for entailment tree construction (Bostrom et al., 2022; Hong et al., 2022). (Ribeiro et al., 2022) also use iterative retrieval-conditioned generation, and (Yang et al., 2022; Tafjord et al., 2022) also use entailment classifiers to filter proof steps. Few tree generation approaches consider this harder scenario of multiple-choice QA, opting instead to focus on the reconstruction task from support facts. (Tafjord et al., 2022) do consider the harder task. They propose a backward chaining QA system that generates entailment trees (not grounded in human-verified facts) using a search driven by internal model belief of factuality, complementary to Nellie’s use of guided and retrieval-conditioned generation to obtain knowledge grounding.
3 Background
[…]
Neural Predicates While most declarative predicates have meaning only in the context of user-defined inference axioms, others can call external neural modules that produce values for their arguments or determine the truth value of the predicate. A popular implementation of this paradigm is DeepProbLog (Manhaeve et al., 2018). We use this feature for defining neural model-invoking predicates. In the above example, we might train a seq2seq to produce other names for a disease $Y$, turning $\textsc{OtherName}(Y^{+},X^{-})$  [^2] In Prolog syntax, ‘ $+$ ’ denotes inputs, ‘ -- ’ outputs. into a neural predicate. This mechanism opens up the capacity to introduce externally-defined object symbols, e.g. seq2seq-generated NL, into the engine’s vocabulary.
Weak Unification In classical backward chaining, a unification operator assigns equivalence to two logical atoms; this requires atoms are of the same arity and have no unequal ground literals in the same argument position. Issues arise when literals are NL sentences, which can be syntactically distinct but semantically equivalent. To handle this, (Weber et al., 2019) propose a weak unification operator, which allows for the unification of any same-arity atoms regardless of conflicting symbols. [^3] Introducing weak unification greatly increases the search runtime, as one might try to unify any two symbols in the vocabulary at every recursive step. This poses a substantial challenge when applying the query grounding algorithm to a search space over NL. They estimate a unification score as the aggregation of pairwise similarity scores using a similarity function $\theta(s_{1},s_{2})\in[0,1]$. The score of the full proof is taken as the minimum of scores across all steps. In this work, we apply a similar aggregation; we say that a query fact $s_{1}$ “weakly unifies” with provenance fact $s_{2}$ with unification score equal to the confidence of an NLI model taking $s_{2}$ as the premise and $s_{1}$ as the hypothesis.
Figure 3: Proposed system framework. An off-the-shelf theorem prover searches for proofs of query $\textsc{Prove}(h)$, where symbol $h$ is an NL hypothesis translated from a QA pair. The prover uses a set of meta-axioms invoking neural retrieval, entailment, and generation predicates to dynamically instantiate inference rules that use the NL factbase.
4 Overview of Approach
Depicted in Figure 3, our proposed framework is comprised of: An external corpus of facts (some $\{f_{1},\dots f_{n}\}$); a module that converts a QA pair to a hypothesis; an off-the-shelf theorem prover; and a suite of meta-axioms that use neural fact retrieval and dynamic rule generation modules to propose, verify, and score inferences. In our experiments, we consider one implementation of this framework that uses the corpus WorldTree (Xie et al., 2020), a set of 9K NL science facts that can interchangeably be considered as rows in 81 $n$-ary relational tables.
Question Conversion Given a multiple-choice question, the system converts each candidate answer into a hypothesis $h$ using a Question to Declarative Sentence model (QA2D; Demszky et al., 2018) (See §C). It then searches for a proof of $h$ against its knowledge base For each alternative, we enumerate $p$ proofs using a time-capped backward chaining search, and then take as the system’s answer the candidate with the overall highest-scoring proof.
4.1 Inference Rule Structure
Our approach uses LMs to dynamically generate inference rules given a hypothesis. The rule structure is strikingly simple, instantiating one of the following meta-level templates:
  1. I.
    Hypothesis $\Leftarrow$ Fact
  2. II.
    Hypothesis $\Leftarrow$ Fact1 $\land$ Fact2
Via template I, the system proves the hypothesis by finding a provenance Fact stored in its knowledge store that entails the hypothesis. Via template II, it enumerates a pair, Fact1 and Fact2, both either stored in the knowledge store or themselves recursively proved, such that the pair in conjunction entail the hypothesis.  [^4] Note that a strict two-premise decomposition can prove expressively powerful. For example, to prove a hypothesis such as those in EntailmentBank (Dalvi et al., 2021) that require a three-premise conjunction $h\Leftarrow f_{1}\land f_{2}\land f_{3}$, Nellie produces instead a recursive set of decompositions $H\Leftarrow f_{1}\land f_{i};f_{i}\Leftarrow f_{2}\land f_{3}$. Template I is given higher search precedence than II, yielding an intuitive high-level procedure: we first look up the hypothesis against the factbase, searching for an entailing fact. If we do not find one, we decomposes the hypothesis into a pair of statements to be proved. Concretely, for an input hypothesis $h$, we define the predicate $\textsc{Prove}(h)$ that serves as the primary goal term. We define the following core meta-rules, which use the neural predicates Retrieve, Entails, and RuleGen. At each step in the backward-chaining search, Nellie’s Prolog engine attempts unify a query with the head of one of these three rules:
  1. 1.
    Fact Unification
    $\textsc{Prove}(h)\Leftarrow\textsc{Retrieve}(h^{+},f^{-})\land\textsc{Entails}(f,h)$
  2. 2.
    Two Premise Rule Generation
    $\textsc{Prove}(h)\Leftarrow\textsc{RuleGen}(h^{+},f_{1}^{-},f_{2}^{-})\ \land$ $\textsc{Entails}([f_{1},f_{2}],h)$
    $\quad\land\ \textsc{Prove}(f_{1})\land\textsc{Prove}(f_{2})$
  3. 3.
    Retrieved First Premise Rule Generation
    $\textsc{Prove}(h)\Leftarrow\textsc{Retrieve}(h^{+},f_{1}^{-})\ \land\ \textsc{RuleGen}(h^{+},f_{1}^{+},f_{2}^{-})$ $\quad\land\textsc{Entails}([f_{1},f_{2}],h)\ \land$ $\textsc{Prove}(f_{2})$
4.2 Unification with Retrieved Facts
For factbase fact $f$, $\textsc{Prove}(f)$ is vacuously true. Rule 1 shows how we prove $\textsc{Prove}(h)$ using retrieval. The predicate Retrieve proposes candidate $f_{i}$’s given $h$ using a FAISS (Johnson et al., 2019)-based nearest neighbor dense retrieval index. [^5] For some situational questions, it is necessary to ground a subquery in evidence from the problem, e.g. “A ball is rolling down a hill.” To handle this, we add the question setup (defined as all but its last sentence) as a “fact” always proposed by Retrieve. We train the retrieval encoder via ranking loss such that the embedding for a hypothesis is maximally similar to its supporting facts. To promote logical coherence and improve the precision of the system, we apply a set of differently-trained neural RTE models as filters $\textsc{Entails}_{j}(\cdot)$ that iteratively rule out $f_{i}$ candidates that are not classified as entailing $h$.
$\textsc{Entails}(f_{i},h)\Leftarrow\bigwedge_{j=1\dots n}\textsc{Entails}_{j}(f_{i},h)$
Implicit in rule 1 is that $\textsc{Prove}(h)$ weakly unifies with some $\textsc{Prove}(f_{i})$; we assign the unification score $\theta(h,f_{i})$ equal to the confidence of one RTE model.
4.3 Dynamic Rule Generator (DRG)
If we do not find an entailing fact for hypothesis $h$, we to decompose $h$ into a pair of entailing premises; we propose candidates using nucleus sampling (Holtzman et al., 2019) from a seq2seq model. The predicate $\textsc{RuleGen}(h,f_{1},f_{2})$ prompts a model trained to generate $h\rightarrow f_{1},f_{2}$ pairs. The space of potential decompositions is very large; to bias the proof search towards those more likely to be grounded in the corpus, we adopt a two-pronged approach, illustrated in inference rules 2 and 3, that proposes a heterogeneous search frontier using different biasing strategies.
Template Conditioned Generation One way in which we might improve our LM-based proposal function is to leverage the high-level structure that supports reasoning in a given domain through template-guided generation, in this case to bias search towards the semi-structure of WorldTree. [^6] While our approach is applicable to a wide array of reasoning problems, the use of WorldTree-specific templates illustrates one way by which we can infuse domain-specific structure into the neural search algorithm. We note that this is a strict departure from the burdensome process of symbolic knowledge engineering; it is instead a simple means by which we can enlist a subject matter expert in order to improve the effectiveness of the engine. As our experiments show, this module improves Nellie’s QA accuracy by a few points, though ablating it yields a similarly strong performance. WorldTree tables correspond to types of facts with similar syntax and semantics that support scientific reasoning. WorldTree QA items are annotated with the facts that support their answers. Each table can be viewed as an n-ary relation of columns whose values are text spans. E.g., the Taxonomic relation has columns <A>, [HYPONYM], <is a kind of>, <SCOPE1>, [HYPERNYM], <for>, [PURPOSE]. Rows include ‘a bird is an animal’ and ‘a seed is a kind of food for birds.’
Accordingly, for the RuleGen predicate in rule 2, half of the $f_{1},f_{2}$ candidates are sampled from the DRG conditioned on $h$ plus a template that cues the model to reflect the syntax of a WorldTree table. We train the DRG to accept any masked infilling template (e.g. <mask>is a kind of <mask>, akin to those used to pretrain recent LMs (Lewis et al., 2019; Raffel et al., 2020)), and propose decompositions whose first fact reflects the template’s syntax. We thus create a $h,t_{1}\rightarrow f_{1},f_{2}$ model. We feed the model templates drawn from WorldTree’s tables, guiding it towards proof steps more likely to be grounded in the factbase. We reuse the same model for non-template-conditioned generation by feeding it an empty template. In practice, we make two generation calls: one samples $m$ free-generated candidates, and a second samples $n$ candidates for each of $n_{t}$ templates (see §C).
Template Selection WorldTree is a diverse corpus, containing tables that are specific to a particular subset of science problems (e.g. the “Predator-Prey” table). As conditioning on dozens of templates can be computationally expensive, we introduce a case-based reasoning (Schank, 1983; Das et al., 2021) approach that selects relevant templates for a given hypothesis. We construct an Okapi-BM25 (Jones et al., 2000) retrieval index over questions from the WorldTree QA training set to obtain the most lexically similar items to a query. At inference time, we select the top-$k$ most similar questions to the query, and take as our template subset the tables of the questions’ annotated facts.
Retrieval Conditioned Generation In rule 3, rather than generate a pair of subqueries, we immediately ground half of the antecedent by choosing as $f_{1}$ a fact retrieved directly from the corpus. We have the DRG force decode $f_{1}$ before generating $f_{2}$ as normal, then recur only on $f_{2}$.
Filters As stochastic sampling from LMs can be a noisy process, some fraction of generated candidates may be invalid: generated facts may be incoherent or not compositionally entail $h$. Accordingly, we introduce a suite of compositional entailment verifiers (Khot et al., 2020; Jhamtani and Clark, 2020) trained on two-premise compositional entailment. We also add a “self-ask” filter, which following (Tafjord et al., 2022) is an LM fine-tuned to assign a statement a truth value. If the confidence is below 0.5 for either an entailment judgement or the ‘self-ask’ belief in $f_{1}$ or $f_{2}$, then we filter the pair. All filters condition on the question setup text. When Nellie uses these rules, the unification score equals the lowest of the scores for $\textsc{Prove}(f_{1})$, for $\textsc{Prove}(f_{2})$, and the confidence of an entailment filter $s_{e}([f_{1},f_{2}]\Rightarrow H)$.
4.4 Proof Search
Given a query, Nellie searches for $t$ seconds to find up to $p$ proofs of depth $d$ or less. We follow (Weber et al., 2019) in pruning search branches whose unification score is guaranteed to fall below the current best, given our monotonic aggregation function $\min(\cdot)$. Found proofs that score under the current best do not count towards $p$. Full pseudocode for the algorithm, which follows a depth-first search with a breadth-first lookahead (Stern et al., 2010) to check for unification of generated subgoals, can be found in §E.
5 Experiments
We train the components of Nellie to be able to answer questions in the Science QA domain. The different neural modules are trained on reformulations of existing datasets for scientific reasoning. Further information can be found in §A.
Our experiments illustrate how the approach exemplified by Nellie provides grounded and logical explanations, performing comparably or better than approaches that do not satisfy these properties. We evaluate models on two multiple-choice QA datasets constructed so that correct answers are supported by facts in the WorldTree corpus: EntailmentBank (Dalvi et al., 2021) is a dataset of entailment trees for declarativized ARC answers, showing how the hypothesis can be derived via compositional entailment hops from WorldTree facts. We recast the test set, initially designed to test tree reconstruction, into QA by retrieving the corresponding multiple-choice ARC questions from which the hypotheses were constructed. WorldTree (Xie et al., 2020) is a subset of the ARC dataset (Clark et al., 2018) annotated with undirected explanation graphs whose nodes are facts from the WorldTree tablestore. We note that WorldTree explanations do not show how facts should combine. There is no guarantee that fully grounded trees exist for these questions using the WorldTree corpus alone. The difficulty of this task is akin to that of a course exam: the teacher (us) provides the student (the model) a long study guide of facts, but expects the student to use these facts via novel composition in order to reason coherently about a problem.
Our task metric is accuracy: whether a generated proof of the correct option outscores any other. [^7] If it produces no proofs, it gives no answer and is wrong. Nellie searches for up to $p$=10 proofs of max depth $d$=5 with a timeout of $t$=180 seconds per option. §B discusses further hyperparameters.
Explanations QA Accuracy (%)
Grounded Logical Ovr Easy Chal
EntailmentBank QA
Nellie (3B) Yes Yes 71.4 76.4 60.4
Entailer-3B No Yes 48.7 52.8 39.6
 (D $=$ 1) 64.9 71.2 50.9
Entailer-11B No Yes 73.2 77.3 64.2
 (D $=$ 1) 71.1 76.4 59.4
WorldTree QA
Nellie (3B) Yes Yes 71.4 75.7 60.9
Entailer-3B No Yes 45.2 50.1 33.3
 (D $=$ 1) 47.7 51.6 38.5
Entailer-11B No Yes 73.2 76.7 64.6
 (D $=$ 1) 74.1 78.7 63.0
PathNet Yes No 43.4
TupleILP Yes No 49.8
ExplanationLP Yes No 62.6
Diff-Explainer Yes No 71.5
Table 1: Nellie performance vs. comparable XQA systems.
Baselines We evaluate Nellie against Entailer (Tafjord et al., 2022), another system that produces entailment tree proofs via backchaining. Entailer stops recurring not when it finds entailing facts from a corpus, but rather when the model believes that a subquery is true with high confidence. Its trees are thus not grounded in verified facts. We reimplemented their algorithm using their T5-11B-based model, reporting their configuration of a max tree depth (D) of 3 and minimum of 1 (i.e. $\geq 1$ decomposition). We also evaluate their ablated setting which recurs exactly once (D=1). This is a baseline that is neither grounded nor particularly interpretable, generating just a pair of statements. To isolate the impact of model size, we also evaluate an Entailer-3B model based on the same T5-3B model as Nellie’s DRG. [^8] We obtained the training data from (Tafjord et al., 2022) and trained our model in consultation with the authors.
We also compare against grounded xQA methods without logical structure (see Fig 2): PathNet (Kundu et al., 2019), which constructs 2-hop chains by linking entities between facts, and three approaches that build graphs from facts using linear programming: TupleILP (Khot et al., 2017), ExplanationLP (Thayaparan et al., 2021), and Diff-Explainer (Thayaparan et al., 2022). [^9] We show numbers reported by (Thayaparan et al., 2021, 2022).
Figure 4: Effect of ablating one or both of rule-conditioned (RCG) and template-conditioned generation (TCG)
Figure 5: (Left) Example Nellie proofs. Top-level queries are decomposed into subqueries via retrieval- or template-conditioned generation. Proof leaves are corpus facts. (Right) Common classes of error causing false statements to be grounded in true ones.
Figure 6: Nellie QA accuracy and proof recall on OBQA with vs. without access to common knowledge statements
5.1 Results
Table 1 shows QA performance. Nellie’s overall (Ovr) accuracy of over 70% matches that of the best-performing grounded baseline, Diff-Explainer, on WorldTree, and is within 2 points of the best logically directed baseline, Entailer-11B. This is notable given Entailer-11B is a much larger model and has no requirement to provide grounded proofs. Nellie outperforms the same-sized Entailer-3B baseline by a margin of 22% on EntailmentBank and 26% on WorldTree under its default max D = 3 configuration, and even outperforms the less explainable D = 1 variant.
Figure 4 shows the results of ablating Nellie’s two conditional generation modules and replacing them with vanilla generation. Removing both template- and retrieval-conditioned generation lowers Nellie’s performance in all circumstances, highlighting the empirical benefit of structured guidance. Ablating TCG (Nellie row vs. (-- TCG) and (-- RCG) vs. (-- Both)) drops performance by 1-1.5 points. Ablating RCG (Nellie vs (-- RCG)) drops it by 3-4.
5.2 Domain Generalization and Knowledge Scaling
While Nellie was trained on datasets centered around WorldTree, we show that it can perform in another domain by altering the knowledge over which it reasons. We consider OpenBookQA (Mihaylov et al., 2018), a dataset that requires reasoning over facts not included in WorldTree; each question is associated with one WorldTree science fact (“metals conduct electricity”), but also one fact from a separate pool of common knowledge (“a suit of armor is made of metal”).
As with a classical expert system, we have designed Nellie to be “improvable merely by making statements to it” (McCarthy, 1959). This means that as we increase its provided knowledge store, we should expect Nellie to reason more effectively. Because WorldTree does not cover all the knowledge required to answer OBQA questions, we test whether Nellie performs better on them when we add the common knowledge to its retrieval index. As the common knowledge annotations are one fact per question and are not designed specifically for entailment, we note that it is unlikely a priori that a fully grounded entailment tree can be created using the available knowledge alone, making this a challenging task for our reasoner.
Figure 7: Distribution of Nellie error classes among 136 incorrect answers on EntailmentBank questions
Figure 6 shows Nellie’s accuracy with and without RCG. We do not use TCG, as the targeted common knowledge is free-form. We find that QA accuracy increases 11-12% with the common knowledge added. The percent of correct statement proved also increases 9-12%. These results also highlight the importance of RCG, as it provides a 5% QA boost and 13% on correct proof rate. These results show that Nellie can be applied out-of-the-box to a dataset requiring reasoning over a different source of free-form NL knowledge.
5.3 Tree Error Analysis
We find that Nellie can produce high quality proofs of correct hypotheses; examples are shown in Figures 5 (left) and §F. However, the system can also generate proofs for incorrect answers that bypass its entailment filters. We can diagnose (and perhaps annotate) error patterns to address in future work. We list a few categories, depicted in Figure 5 (right). A common pattern of error is Hypernym Property Inheritance Errors, in which the model assumes that a member of a taxonomic class has a property of their hypernym, but the property is not universal. A colloquial example is inferring penguins can fly from penguins are birds & birds can fly. We also observe an amount of Errors from Semantic Drift (Khashabi et al., 2019) between inference hops, culminating in RTE model false positives. E.g., in block $2$ of Fig 5 (right), which object loses temperature during heat transfer changes between hops. Figure 7 shows the distribution of these errors on a set of 136 incorrect answers from EntailmentBank. The predominant error case is a failure to find any proof of the correct option.
6 Conclusion
We have proposed a reimagined version of a classical expert system that relies on the inferential power of LLMs rather than handcrafted rules. Nellie has the skeleton of a symbolic theorem prover, but the provided rules invoke neural predicates that allow for systematic reasoning over NL statements. In order to dynamically generate inferences, we introduce two mechanisms for knowledge-guided generation: conditioning decompositions on retrieval or model-selected inference templates. Our search algorithm undergirds Nellie, the first explainable reasoning system that performs end-to-end QA by searching for grounded entailment trees. We find that Nellie equals or exceeds the performance of comparable XQA systems, while providing explanations with the simultaneous guarantees of logical directionality and groundedness in human-provided text. This work thus suggests a new way to jointly reap the benefits of both modern neural methods and traditional reasoning.
Appendix A Model Details and Training
We train the components of Nellie to be able to answer questions in the Science QA domain. The different neural modules are trained on reformulations of existing datasets for scientific reasoning. We use the following sources: WorldTree Explanations [Xie et al., 2020] and EntailmentBank [Dalvi et al., 2021] are used as described in the paper. eQASC [Jhamtani and Clark, 2020] and SciTaiL [Khot et al., 2018] are datasets of two- and one-premise entailment questions.
QA2D Our QA2D model is a T5-11B [Raffel et al., 2020] fine-tuned by [Tafjord et al., 2022] on the original QA2D dataset [Demszky et al., 2018].
Retrieval Module The dense retrieval model is a SentenceTransformers Siamese BERT-Network encoder [Reimers and Gurevych, 2019] pretrained on the MS MARCO IR corpus [Nguyen et al., 2016], which we fine-tune via ranking loss to maximize the cosine similarity between a hypothesis and its support facts. We gather $h,f_{1}$ from the WorldTree, EntailmentBank, and eQASC training sets, then sample random negatives during training.
Dynamic Rule Generator The DRG is a T5-3B model [Raffel et al., 2020] fine-tuned on $h\rightarrow[f_{1},f_{2}]$ pairs drawn from eQASC, which are numerous (26K) but noisy, and the binary composition steps from the EntailmentBank training set, which are few (3.8K) but ultimately rooted in facts from WorldTree. To make the DRG able to condition on templates, we adopt the span masking strategy from [Raffel et al., 2020] to create a randomly masked version $\hat{f}_{1}$ of each $f_{1}$ and then train the model on $h,\hat{f}_{1}\rightarrow f_{1},f_{2}$ pairs. For each original triplet, we create one training example with a template, and one in which $\hat{f}_{1}$ is an empty <mask>, giving the model the dual capacity for non-template-generation.
Entailment Filters We find that a mixture of strategies helps prioritize precision over recall when filtering candidates. For 1-premise filtering for unification and 2-premise filtering for rule generation, we fine-tune a separate pair of models consisting of a SentenceTransformer cross-encoder classifier and a T5 seq2seq. The former is trained via classification loss, the latter via sequence cross entropy loss. The 1-premise filters are trained on the SciTail dataset, the 2-premise filters on eQASC and EntailmentBank. For 2-premise entailment, we additionally add a “self-ask” filter, which following [Tafjord et al., 2022] is a T5 model fine-tuned to produce its 0-to-1 confidence that a statement is true. If the confidence is below 0.5 for either $f_{1}$ or $f_{2}$, we filter the pair. As in [Tafjord et al., 2022], we train our DRG seq2seq in a multitask fashion so that it serves as both the rule generator as well as the “self-ask” filter and (half of) the 2-premise entailment filter.
A.1 Training
We use the default training parameters from the SentenceTransformers library [^10] https://tinyurl.com/mvhbmfcc in order to train CrossEncoder-based classifier and Bi-encoder retriever models for 3 epochs. We correspondingly use the default fine-tuning parameters from the HuggingFace library in order to train our T5-based generation and classification models.
Appendix B Hyperparameters
[…]
We used Optuna [Akiba et al., 2019] to find the number of facts for retrieval-conditioned generation, the $p$ value for nucleus sampling, the entailment threshold, and the number of candidates generated at each recursion step. We used as our objective the overall QA accuracy on the EntailmentBank dev set. We ran it for about 300 trials, and found that the most important parameters were the entailment threshold (too low = too many bad proofs, too high = no proofs) and the nucleus $p$ (less than 0.8 reduced QA performance by 10-15. Experiments were run on 24GB Quadro RTX 6000s.
WorldTree Relation Template
INSTANCES <mask>is <mask>
ACTION <mask>to <mask>
KINDOF <mask>is a kind of <mask>
PROP-THINGS <mask>are <mask>
ACTION <mask>for <mask>
AFFORDANCES <mask>can <mask>
IFTHEN if <mask>then <mask>
PROP-THINGS <mask>has <mask>
CAUSE <mask>causes <mask>
MADEOF <mask>made of <mask>
PROP-THINGS <mask>have <mask>
REQUIRES <mask>requires <mask>
ACTION <mask>is when <mask>
USEDFOR <mask>used to <mask>
PARTOF <mask>a part of <mask>
CHANGE <mask>changes <mask>
USEDFOR <mask>used for <mask>
ATTRIBUTE-VALUE-RANGE <mask>means <mask>
AFFECT <mask>has <mask>impact on <mask>
PROP-ANIMAL-ATTRIB <mask>is a <mask>animal
SOURCEOF <mask>is a source of <mask>
CONTAINS <mask>contains <mask>
COUPLEDRELATIONSHIP as <mask>will <mask>
CHANGE-VEC <mask>decreases <mask>
COMPARISON <mask>than <mask>
ACTION <mask>is when <mask>to <mask>
CAUSE <mask>can cause <mask>
LOCATIONS <mask>found <mask>
COMPARISON <mask>more <mask>
PROP-INHERITEDLEARNED <mask>is <mask>characteristic
DURING <mask>during <mask>
CHANGE-VEC <mask>increases <mask>
AFFORDANCES <mask>can <mask>by <mask>
EXAMPLES an example of <mask>is <mask>
SOURCEOF <mask>produces <mask>
CONTAINS <mask>contain <mask>
CHANGE <mask>converts <mask>
COMPARISON <mask>is <mask>than <mask>
FORMEDBY <mask>formed by <mask>
CONSUMERS-EATING <mask>eat <mask>
SOURCEOF <mask>provides <mask>
PROP-RESOURCES-RENEWABLE <mask>is <mask>resource
ATTRIBUTE-VALUE-RANGE <mask>means <mask>of <mask>
COMPARISON <mask>less <mask>
COUPLEDRELATIONSHIP as <mask>decreases <mask>will <mask>
PROP-CONDUCTIVITY <mask>is <mask>conductor
PROP-GENERIC <mask>is a property of <mask>
HABITAT <mask>live in <mask>
MEASUREMENTS <mask>is a measure of <mask>
COUPLEDRELATIONSHIP as <mask>increases <mask>will <mask>
Figure 8: Partial list of WorldTree templates used for guided generation, sorted by frequency of occurrence in EntailmentBank trees (full list omitted for space).
Appendix C QA2D Conversion
The following is an example conversion by our QA2D model of a multiple-choice question into a set of candidate hypotheses to prove.
Q: Ethanol is an alternative fuel made from corn. What is one of the unfavorable effects of using ethanol as a fuel?
  1. (A)
    decreased cost of fuel production
  2. (B)
    decrease in farm land available for food production
  3. (C)
    increase in the consumption of fossil fuels
  4. (D)
    increased carbon footprint from driving automobiles
Converted to the following hypotheses:
  1. (A)
    Using ethanol has an unfavorable effect on the cost of fuel.
  2. (B)
    Using ethanol as fuel decreases farm land available for food production.
  3. (C)
    Using ethanol as fuel increases in the consumption of fossil fuels.
  4. (D)
    The use of ethanol as an alternative fuel can lead to increased carbon footprints.
Appendix D WorldTree Templates
table 8 depicts the infilling templates gathered from WorldTree and used for template-guided rule generation. We manually annotated this list. The full set of templates is of size $150$.
Appendix E Search Algorithm
Algorithm 1 depicts Nellie’s search algorithm, the logic of which is implemented in DeepProbLog [Manhaeve et al., 2018]. The search is parametrized by
  1. 1.
    A maximum number of proofs $m$ at which to cut off searching. We set this to $10$ for top-level queries and $2$ for recursive subqueries.
  2. 2.
    Candidate generation rates $n_{v}$ for vanilla nucleus-sampled decompositions, $n_{t}$ for template-conditioned decompositions, and $n_{r}$ for retrieval-conditioned generations. We set these to $200$, $250$ and $200$, respectively. [^11] Due to batching, these correspond to 1, 2, and 1 total calls to HuggingFace’s generate function. Upon removing exact match duplicates, a call to RuleGen produces about $400$ candidates.
  3. 3.
    The number of support facts $n_{f}$ to retrieve at each call to $\textsc{Retrieve}_{K}$, which we set to 15.
  4. 4.
    Entailment scoring module $s_{e}(\cdot)$, which is a separate NLI cross-encoder model for single- and double-premise entailments.
1:procedure Prove($H$, $m$) $\rightarrow$ proofs $tree^{(1)}(H),\dots tree^{(m)}(H)$ scores $s^{(1)}(H),\dots s^{(m)}{H}$:
2:   i. check weak unification against $K$:
3:       entailing premises $P={\color[rgb]{0,0,1}\textsc{PremiseLookup}}(H)$
4:       if $P$ nonempty
[…]
6:   ii. generate candidate decompositions: // If $H$ does not unify against $K$, decompose it
7:       decompositions $D={\color[rgb]{0,0,1}\textsc{RuleGen}}(H)$
8:   iii. check whether any decompositions pairs both unify: // this step amortizes premise lookups
9:       proofs = []; scores = []
10:       forall $(p_{1},p_{2})\in D$ do
11:               if $p_{1}\in K$  and  $\exists p_{2}^{\prime}\in{\color[rgb]{0,0,1}\textsc{PremiseLookup}}(p_{2})$ // if $p_{1}$ already grounded, its subtree is leaf
12:               then proofs += ($(\{p_{1},(p_{2}^{\prime}\vdash{}p_{2})\}\vdash{}H)$
13:                   scores += $\min({\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{2}^{\prime}\vdash{}p_{2}),{\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{1}\land p_{2}\vdash{}H))$ // tree score is min of scores used to construct it
14:               elif $\exists p_{1}^{\prime}\in{\color[rgb]{0,0,1}\textsc{PremiseLookup}}(p_{1})$  and  $\exists p_{2}^{\prime}\in{\color[rgb]{0,0,1}\textsc{PremiseLookup}}(p_{2})$
15:               then proofs += ($(\{(p_{1}^{\prime}\vdash{}p_{1}),(p_{2}^{\prime}\vdash{}p_{2})\}\vdash{}H)$
16:                   scores += $\min({\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{1}^{\prime}\vdash{}p_{1}),{\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{2}^{\prime}\vdash{}p_{2}),{\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{1}\land p_{2}\vdash{}H))$
17:       if proofs not empty
18:       then return proofs, scores
19:   iv. recur on generated premises: // if still no proof, need to recur
20:       proofs = []; scores = []
21:       forall $(p_{1},p_{2})\in D$ do
22:               if ${\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{1}\land p_{2}\vdash{}H)>\max(\text{scores})$  and  $\text{len}(\text{proofs})<m$ // only recur if chance of better score
23:               then $(\text{proof}_{p_{1}},S_{p_{1}})={\color[rgb]{0,0,1}\textsc{Prove}}(p_{1},1)$
24:                   $(\text{proof}_{p_{2}},S_{p_{2}})={\color[rgb]{0,0,1}\textsc{Prove}}(p_{2},1)$
25:                   recursive score $S_{r}=\min(S_{p_{1}},S_{p_{2}},{\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{1}\land p_{2}\vdash{}H))$
26:                   if $S_{r}>\max(\text{scores})$
27:                   then proofs += ($(\{\text{proof}_{p_{1}},\text{proof}_{p_{2}}\}\vdash{}H)$
28:                      scores += $S_{r}$
29:       return proofs, scores
30:
31:procedure PremiseLookup(H) $\rightarrow$ premises $P$ s.t. $\forall p\in P,p\vdash{}H$ // try to ground H in K via entailment
32:       candidate premises $P={\color[rgb]{1,.5,0}\textsc{Retrieve}_{K}}(H,n_{f})$ // retrieve candidate premises s.t. $p\vdash{}H$
[…]
34:               $P=f(P,H)$ // only keep premises that pass all filters
35:       return $P$
36:
37:procedure RuleGen(H) $\rightarrow$ decompositions $D=(p_{1}^{1},p_{2}^{1}),\dots(p_{1}^{c},p_{2}^{c})$ // find candidate pairs s.t. $p_{1}\land p_{2}\vdash{}H$
38:   D = []
39:   $D$ += ${\color[rgb]{1,.5,0}\textsc{Generate}}(H,n=n_{v})$
40:   WTTemplates = CaseBasedRetrieveTemplate(H)
41:   outputs per template $o_{t}=n_{t}/\text{len}({\color[rgb]{0.56,1,0.26}\textsc{WTTemplates}})$
42:   forall $t\in{\color[rgb]{0.56,1,0.26}\textsc{WTTemplates}}$ do // generate candidates for each of 50 templates in WTTemplates
43:       $D$ += ${\color[rgb]{1,.5,0}\textsc{Generate}}([H\ t],n=o_{t})$ // T5 model accepts optional $t$ appended to $H$
44:   retrieved first premises $P_{1}={\color[rgb]{1,.5,0}\textsc{Retrieve}_{K}}(H,n_{f})$
45:   outputs per premise $o_{r}=n_{r}/\text{len}(P_{1})$
46:   forall $p_{K}\in P_{1}$ do // generate candidates with retrieved support facts as first premise
47:         $D$ += ${\color[rgb]{1,.5,0}\textsc{Generate}}(H,\text{prefix}=p_{K},n=o_{r})$ // Force decode $p_{K}$ then generate a $p_{2}$
[…]
49:               $D=f(D,H)$ // only keep premise pairs that pass all filters
50:         return $D$
Algorithm 1 Nellie’s backchaining algorithm, coded in Prolog, for searching for proof list $tree^{(1)}(H),\dots tree^{(m)}(H)$ with scores $s^{(1)}(H),\dots s^{(m)}(H)$ for a hypothesis $H$ against factbase $K$. The rule generator produces $n_{v}$ candidates via vanilla nucleus sampling, $n_{t}$ candidates via sampling from templates, and $n_{r}$ candidates whose first fact is retrieved from K. All Generate, Retrieve, and entailment scorer $\textsc{Score}_{e}$ calls in forall loops are batched. All module outputs are cached. OnePremiseFilters and TwoPremiseFilters  are ordered lists of neural entailment models.
Appendix F Example Proofs
Figure 9 illustrates more proofs generated by Nellie for questions from EntailmentBank and WorldTree. Blue-outlined nodes are hypotheses generated by the QA2D model for the correct answer highted in green. Grey nodes are generated subqueries; tan nodes are leaves from Nellie’s knowledge corpus. The green node is a question setup text that grounds a contextual subquery via entailment.
Question
Nellie Proof
Cheetahs have come close to extinction due to hunting, drought, and disease. There is now very little genetic variation in cheetah populations. Which of the following is a result of the limited genetic variation in the current cheetah populations compared to earlier cheetah populations with more variation? (A) Cheetahs in current populations are more resistant to new diseases. (B) The survival rate of young cheetahs is increased in current populations. (C) Cheetahs in current populations are less able to interbreed with other species. (D) The current cheetah populations are less likely to be able to adapt to environmental changes.
In a city near the ocean, fog often forms on summer mornings. Which of the following statements best explains how this fog forms? (A) Ocean water evaporates and then condenses in the air. (B) Crashing waves spray tiny drops of ocean water into the air. (C) Water runoff moves toward the ocean and collects near the shore. (D) Rain clouds move in from the ocean and evaporate as they reach the shore.
The snowshoe hare sheds its fur twice a year. In the summer, the fur of the hare is brown. In the winter, the fur is white. Which of these statements best explains the advantage of shedding fur? (A) Shedding fur keeps the hare clean. (B) Shedding fur helps the hare move quickly. (C) Shedding fur keeps the hare’s home warm. (D) Shedding fur helps the hare blend into its habitat.
Scientists use large optical telescopes to obtain information about the planets in the solar system. What wavelengths of electromagnetic radiation provide this information? (A) gamma radiation (B) infrared radiation (C) radio waves (D) visible light
Figure 9: Example multiple-choice questions with NELLIE’s answer and corresponding proof.

### v5

@@
 Abstract
-Our goal is a modern approach to answering questions via systematic reasoning where answers are supported by human interpretable proof trees grounded in an NL corpus of authoritative facts.
+Our goal is to develop a modern approach to answering questions via systematic reasoning where answers are supported by human interpretable proof trees grounded in an NL corpus of facts.
 Such a system would help alleviate the challenges of interpretability and hallucination with modern LMs, and the lack of grounding of current explanation methods (e.g., Chain-of-Thought).
@@
 Our implementation, Nellie, is the first system to demonstrate fully interpretable, end-to-end grounded QA as entailment tree proof search, going beyond earlier work explaining known-to-be-true facts from text.
-In experiments, Nellie outperforms a similar-sized state-of-the-art reasoner (Tafjord et al., 2022) while producing knowledge-grounded explanations.
+In experiments, Nellie outperforms a similar-sized state-of-the-art reasoner while producing knowledge-grounded explanations.
 We also find Nellie can exploit both semi-structured and NL text corpora to guide reasoning.
@@
 This general opacity is a growing impediment to widespread use of LLMs, e.g., in critical applications such as medicine, law, and hiring decisions, where interpretability and trust are paramount.
-While there has been rapid progress on having LLMs generate systematic explanations for their answers, e.g., Chain-of-Thought (Wei et al., 2022), Entailer (Tafjord et al., 2022), or Maieutic Prompting (Jung et al., 2022), such explanations are not grounded in external facts and may include hallucinations (Ji et al., 2022).
+While there has been rapid progress in having LLMs generate systematic explanations for their answers, e.g., Chain-of-Thought Wei et al.
+(2022), Entailer Tafjord et al.
+(2022), or Maieutic Prompting Jung et al.
+(2022), such explanations are not grounded in external facts and may include hallucinations Ji et al.
+(2022).
 Rather, what would be desirable - and what this work pursues - is a system that systematically reasons over authoritative text: to support answers with human interpretable proof trees grounded in the text, while not requiring translation to an entirely symbolic formalism.
 Figure 1: Given a query, Nellie performs a neuro-symbolic backward chaining search for proof trees whose leaves are grounded in a corpus of facts.
-First, it generates candidate decomposition rules guided by both retrieved facts and templates.
-Then, it recursively tries to prove the rules’ conditions via entailment from the corpus or further decomposition.
-Our approach is to revisit the behavior of expert systems (Jackson, 1986; Metaxiotis et al., 2002).
+It generates candidate decomposition rules guided by retrieved facts or templates.
+Then, it recursively tries to prove rule conditions via entailment from the corpus or further decomposition.
+Our approach is to revisit the behavior of expert systems Jackson (1986); Metaxiotis et al.
+(2002).
 Expert systems are appealing for their explainable behavior: decisions are made by constructing a well-formed symbolic proof from explicit, formally represented knowledge authored by a knowledge engineer in consultation with a domain expert.
-However, as expert systems are known to be both expensive and brittle (Musen and Van der Lei, 1988), the AI community has turned to neuro-symbolic mechanisms that use large language models (LLMs) to reason over natural language (NL).
-Reasoning over NL does not require a symbolic formalism, and allows the use of inferentially powerful LLMs, but also loses the systematic, verifiable proofs that expert systems produced.
+However, as expert systems are known to be both expensive and brittle Musen and Van der Lei (1988), the AI community has turned to neuro-symbolic mechanisms that use large language models to reason over natural language (NL).
+Reasoning over NL does not require a symbolic formalism and allows the use of inferentially powerful LLMs, but also loses the systematic, verifiable proofs that expert systems produced.
 This motivates our pursuit of a new way to jointly reap the benefits of both modern neural methods and traditional reasoning.
@@
 As illustrated in Figure 2, various eXplainable QA (XQA) methods accomplish 2 of these criteria, but not all 3.
-Some methods, like Chain-of-Thought, extract inference chains or logically structured graphs fully from a LLM without grounding in verified knowledge; others, like ExplanationLP (Thayaparan et al., 2021), compose graphs of grounded facts, but do not show logical direction; finally, LAMBADA (Kazemi et al., 2023) achieves both direction and grounding, but is limited to simple domains with provided NL Horn rules and small sets of facts (1-2 dozen).
-In contrast, we desire a system that handles a much larger corpus (10K statements) and doesn’t need handcrafted rules.
+Some methods, like Chain-of-Thought, generate inference chains or logically structured graphs from an LLM without grounding in verified knowledge.
+Others, like ExplanationLP Thayaparan et al.
+(2021), compose graphs of grounded facts, but do not show logical direction.
+LAMBADA Kazemi et al.
+(2023) achieves both direction and grounding but is limited to simple domains with provided NL Horn rules and sets of 1-2 dozen facts.
+In contrast, we desire a system that handles a larger corpus (10K statements) and doesn’t need handcrafted rules.
 Figure 2: Comparison of approaches to neural XQA.
@@
 The “symbols” that our neuro-symbolic engine reasons over are free-form NL sentences.
-Rather than require a knowledge engineer to carefully write hundreds of inference rules as in the classical setting, Nellie employs an LLM as a dynamic rule generator  (DRG; Kalyanpur et al., 2021) to generate (rather than retrieve) candidate rules that decompose a hypothesis into subqueries that, if themselves proved recursively, will prove that hypothesis via composition (Figure 1).
+Rather than require a knowledge engineer to carefully write hundreds of inference rules as in the classical setting, Nellie employs an LLM as a dynamic rule generator [DRG; Kalyanpur et al.
+(2021)] to generate (rather than retrieve) candidate rules that decompose a hypothesis into subqueries that, if themselves proved recursively, will prove that hypothesis via composition (Figure 1).
 In this way, Nellie can answer the question posed by the classical expert system: “Does this statement systematically follow from my knowledge, or not?”.
 Nellie is built upon a backward chaining symbolic theorem prover written in Prolog, implemented using a few meta-rules specifying how inference should proceed.
-These include checking for entailment of a hypothesis against a retrieved fact, or decomposing it into a conjunction of subqueries to recursively prove.
-To treat NL sentences as if they were symbols in a purely symbolic search algorithm, we use an NLI-based form of weak unification (Sessa, 2002; Weber et al., 2019) between a queried statement and a corpus fact.
-This leads to interpretable proofs similar to the compositional entailment trees of prior work (e.g., EntailmentBank (Dalvi et al., 2021)), while tackling the additional challenge of multiple-choice QA.
+These include checking for entailment of a hypothesis against a retrieved fact or decomposing it into a conjunction of subqueries to recursively prove.
+To treat NL sentences as if they were symbols in a purely symbolic search algorithm, we use a natural language inference-based form of weak unification Sessa (2002); Weber et al.
+(2019) between the hypothesis and a corpus fact.
+This produces interpretable proofs similar to the compositional entailment trees of prior work (e.g., EntailmentBank Dalvi et al.
+(2021)) while tackling the additional challenge of QA.
 Nellie is designed to search across hundreds of trees whose leaves come from one of two types of knowledge: (a) a corpus of semi-structured text statements, such as NL renditions of database entries, or (b) a corpus of free-form NL sentences.
-Many correct proofs might exist for a given hypothesis, but much fewer will be fully groundable in the provided corpus, making the search harder than for ungrounded alternatives like Entailer (Tafjord et al., 2022) or COMET-Dynagen (Bosselut et al., 2021).
+Many correct proofs might exist for a given hypothesis, but much fewer will be fully groundable in the provided corpus, making the search harder than for ungrounded alternatives like Entailer Tafjord et al.
+(2022) or COMET-Dynagen Bosselut et al.
+(2021).
 To improve grounding, we introduce two guiding methods to boost the likelihood of generating rules whose conditions match against the available text: (1) For applications where a semi-structured corpus is available, Nellie leverages this structure via templates to help steer rule generation towards syntax that is more likely to match corpus entries.
@@
 Ablation experiments show these together and individually improve Nellie’s reasoning.
-Nellie expands upon work explaining known-to-be-true facts from text (SCSEARCH; Bostrom et al., 2022),(IRGR; Ribeiro et al., 2022) to perform QA, including searching for and comparing trees for conflicting options (both true and false).
-Experiments find Nellie outperforms a similar-sized SOTA reasoner (Entailer) while producing compositional trees showing how all decisions (top-level and intermediate) are grounded in corpus facts.
+Nellie expands upon methods that ground known-to-be-true hypotheses to provided facts [SCSEARCH; Bostrom et al.
+(2022)], extending the paradigm to perform QA.
+This involves searching for and comparing trees for conflicting answer options.
+Experiments find Nellie outperforms a similar-sized state-of-the-art reasoner, Entailer, while producing compositional trees showing how decisions are grounded in corpus facts.
 Our contributions are thus:
@@
   2. 2.
-    An implementation, Nellie, [^1] Code will be released at https://github.com/JHU-CLSP/NELLIE that outperforms a similar-sized SOTA reasoner (Entailer-3B) while producing grounded trees.
+    An implementation, Nellie, [^1] Code found at https://github.com/JHU-CLSP/NELLIE.
+that outperforms a similar-sized SOTA reasoner (Entailer-3B) while producing grounded trees.
 To our knowledge, this is the first system to perform grounded XQA as NL entailment tree search.
 2 Related Work
-Theorem Proving over Language A long-standing approach to reasoning over NL is to project it into a symbolic form, such as for QA (Green et al., 1961; Zelle and Mooney, 1996) or entailment (Bos and Markert, 2005).
-Provided the translation from NL to symbolic representation is accurate, one can leverage fast and scalable solutions for discrete symbolic inference (Riazanov and Voronkov, 2002; Kautz et al., 1992; Kautz and Selman, 1999).
-However, reliably translating broad-domain NL into an adequately expressive formalism for reasoning is a challenge (Schubert, 2015).
-Recent work explores methods of handling NL without needing to parse it to another formalism.
-Variants of neural theorem provers (NTPs; Rocktäschel and Riedel, 2017) such as NLProlog (Weber et al., 2019) reconcile NL facts with symbolic reasoning by learning embeddings for both the facts and the symbols in a theorem prover, then using weak unification to backward chain.
-(Kalyanpur et al., 2021) inject neural reasoning into a Boxer/Prolog-based symbolic reasoner via a special LM-calling predicate.
-(Arabshahi et al., 2021b) combine handwritten symbolic rules with neural symbol embeddings in order to classify conversational intents.
-Other work explores using LMs to emulate stepwise (Tafjord et al., 2021; Kazemi et al., 2023) or end-to-end (Clark et al., 2021; Picco et al., 2021) logical reasoning over small rulebases converted to NL.
-Systematic Explanation Generation Recent works have used LMs to generate NL reasoning graphs in support of an answer.
-“Chain-of-Thought” prompting (Wei et al., 2022; Kojima et al., 2022), elicits free-form inference hops from the LM before it generates an answer.
-Other text graph generation methods connect model-generated statements via common sense relations (Bosselut et al., 2021; Arabshahi et al., 2021a) or for/against influence (Madaan et al., 2021; Jung et al., 2022), though these are not knowledge-grounded and don’t address entailment (see Figure 2).
-The EntailmentBank dataset (Dalvi et al., 2021) has driven research towards the construction of explanation trees, challenging models to produce stepwise entailment proofs of a statement using a set of provided support facts.
-This direction builds upon works on explainable reasoning that build graphs from KB-retrieved sets of support statements, but stop short of showing their role in logical entailment (Pan et al., 2019; Jansen and Ustalov, 2019; Yadav et al., 2019; Valentino et al., 2022).
-Components of our framework are related to concurrent approaches for entailment tree construction (Bostrom et al., 2022; Hong et al., 2022).
-(Ribeiro et al., 2022) also use iterative retrieval-conditioned generation, and (Yang et al., 2022; Tafjord et al., 2022) also use entailment classifiers to filter proof steps.
-Few tree generation approaches consider this harder scenario of multiple-choice QA, opting instead to focus on the reconstruction task from support facts.
-(Tafjord et al., 2022) do consider the harder task.
-They propose a backward chaining QA system that generates entailment trees (not grounded in human-verified facts) using a search driven by internal model belief of factuality, complementary to Nellie’s use of guided and retrieval-conditioned generation to obtain knowledge grounding.
+Theorem Proving over Language
+A long-standing approach to reasoning over NL is to project it into a symbolic form, such as for QA Green et al.
+(1961); Zelle and Mooney (1996) or entailment Bos and Markert (2005).
+Provided the translation from NL to symbolic representation is accurate, one can leverage fast and scalable solutions for discrete symbolic inference Riazanov and Voronkov (2002); Kautz et al.
+(1992); Kautz and Selman (1999).
+However, reliably translating broad-domain NL into an adequately expressive formalism for reasoning is a challenge Schubert (2015), though some have found success using LLMs to perform this semantic parsing task Wong et al.
+(2023); Lyu et al.
+(2023); Olausson et al.
+(2023); Pan et al.
+(2023); Ye et al.
+(2023).
+Recent work explores methods of handling NL without parsing it to another formalism.
+Some use LMs to generate proof steps in mathematical theorem proving Polu and Sutskever (2020); Welleck et al.
+(2022).
+Variants of neural theorem provers [NTPs; Rocktäschel and Riedel (2017)] such as NLProlog Weber et al.
+(2019) reconcile NL facts with symbolic reasoning by learning embeddings for the facts and symbols in a theory, then using weak unification to backward chain.
+Kalyanpur et al.
+(2021) inject neural reasoning into a Boxer/Prolog-based symbolic reasoner via a special LM-calling predicate.
+Arabshahi et al.
+(2021b) combine handwritten symbolic rules with neural symbol embeddings to classify conversational intents.
+Other work explores using LMs to emulate stepwise Tafjord et al.
+(2021); Kazemi et al.
+(2023) or end-to-end Clark et al.
+(2021); Picco et al.
+(2021) logical reasoning over small rulebases converted to NL.
+These approaches require both user-provided if/then rules to operate, while Nellie requires only facts and is thus applicable to domains in which rules are not available.
+Modular Reasoning over NL
+Nellie’s systematic reasoning is related to approaches that decompose problems into sequences of modular operations.
+Gupta et al.
+(2020) introduce a neural module network [NMN; Andreas et al.
+(2016)] for QA with modules for span extraction and arithmetic operations.
+Khot et al.
+(2021) introduce another NMN that decomposes questions into simpler ones answerable by existing models.
+These are part of a broader class of work decomposing multi-step reasoning using reasoning modules Khot et al.
+(2022); Saha et al.
+(2023a, b).
+Systematic Explanation Generation
+Recent works have used LMs to generate NL reasoning graphs in support of an answer.
+“Chain-of-Thought” prompting Wei et al.
+(2022); Kojima et al.
+(2022), elicits free-form inference hops from the LM before it generates an answer.
+Other text graph generation methods connect model-generated statements via common sense relations Bosselut et al.
+(2021); Arabshahi et al.
+(2021a) or for/against influence Madaan et al.
+(2021); Jung et al.
+(2022), though these are not knowledge-grounded and don’t address entailment (see Figure 2).
+The EntailmentBank dataset Dalvi et al.
+(2021) has driven research towards the construction of explanation trees, challenging models to produce stepwise entailment proofs of a statement using a set of provided support facts.
+This direction builds upon works on explainable reasoning that build graphs from KB-retrieved sets of support statements, but stop short of showing their role in logical entailment Pan et al.
+(2019); Jansen and Ustalov (2019); Yadav et al.
+(2019); Valentino et al.
+(2022).
+Components of our framework are related to concurrent approaches for entailment tree construction Bostrom et al.
+(2022); Hong et al.
+(2022); Sprague et al.
+(2022).
+Ribeiro et al.
+(2022) also use iterative retrieval-conditioned generation, and Yang et al.
+(2022) also use entailment classifiers to filter proof steps.
+None of the above tree generation approaches consider this harder scenario of multiple-choice QA, opting instead to focus on the reconstruction task from support facts.
+Tafjord et al.
+(2022) do consider the harder task.
+They propose a backward chaining QA system, Entailer, that generates entailment trees (not grounded in human-verified facts) using a search grounded to the model’s internal beliefs, complementary to Nellie’s use of guided and retrieval-conditioned generation to obtain knowledge grounding.
+As with Nellie5.2), Entailer benefits from humans adding more useful statements to its available knowledge Dalvi Mishra et al.
+(2022).
+The difference between systems is highlighted by the number of generations (i.e.
+search nodes) considered by the two algorithms: while Entailer considers at most a couple dozen hypothesis decompositions, Nellie must consider hundreds or thousands to find one that is fully grounded.
+Faithful Complex LLM Reasoning
+Of the growing body of work on LLM-based “Chain-of-X” reasoning methods Xia et al.
+(2024), a portion considers ways to improve faithfulness to underlying knowledge and reduce LLM hallucinations (see Lyu et al.
+(2024) for an overview of faithful explanation methods).
+Nellie contributes to this space by providing a faithful-by-design algorithm that reasons based on logical hypothesis grounding.
+Other recent methods include Rethinking with Retrieval He et al.
+(2022), which reranks reasoning chains using a faithfulness score based on retrieved knowledge, and the forward-chaining Faithful Reasoning algorithm Creswell and Shanahan (2022), which follows “a beam search over reasoning traces” to answer a question by iteratively adding deductive inferences to a context of starting facts.
+Theirs is a different style search to our backward-chaining Nellie, not relying upon NLI for verifying logical connectedness and not scaling to larger knowledge bases.
 3 Background
 […]
-Neural Predicates While most declarative predicates have meaning only in the context of user-defined inference axioms, others can call external neural modules that produce values for their arguments or determine the truth value of the predicate.
-A popular implementation of this paradigm is DeepProbLog (Manhaeve et al., 2018).
-We use this feature for defining neural model-invoking predicates.
-In the above example, we might train a seq2seq to produce other names for a disease $Y$, turning $\textsc{OtherName}(Y^{+},X^{-})$  [^2] In Prolog syntax, ‘ $+$ ’ denotes inputs, ‘ -- ’ outputs.
+Figure 3: Proposed system framework.
+An off-the-shelf theorem prover searches for proofs of query $\textsc{Prove}(h)$, where symbol $h$ is an NL hypothesis translated from a QA pair.
+The prover uses a set of meta-axioms invoking neural retrieval, entailment, and generation predicates to dynamically instantiate inference rules that use the NL factbase.
+Neural Predicates
+While most declarative predicates have meaning only in the context of user-defined inference axioms, others can call external neural modules that produce values for their arguments or determine the truth value of the predicate.
+A popular implementation of this paradigm is DeepProbLog Manhaeve et al.
+(2018), which we use to define LM-invoking predicates.
+In the above example, we might train a sequence-to-sequence (seq2seq) model to produce other names for a disease $Y$, turning $\textsc{OtherName}(Y^{+},X^{-})$  [^2] In Prolog syntax, ‘ $+$ ’ denotes inputs, ‘ -- ’ outputs.
 into a neural predicate.
-This mechanism opens up the capacity to introduce externally-defined object symbols, e.g.
+This mechanism creates the ability to introduce externally defined object symbols, e.g.
 seq2seq-generated NL, into the engine’s vocabulary.
-Weak Unification In classical backward chaining, a unification operator assigns equivalence to two logical atoms; this requires atoms are of the same arity and have no unequal ground literals in the same argument position.
+Weak Unification
+In classical backward chaining, a unification operator assigns equivalence to two logical atoms; this requires atoms to have the same arity and have no unequal ground literals in the same argument position.
 Issues arise when literals are NL sentences, which can be syntactically distinct but semantically equivalent.
-To handle this, (Weber et al., 2019) propose a weak unification operator, which allows for the unification of any same-arity atoms regardless of conflicting symbols.
+To handle this, Weber et al.
+(2019) propose a weak unification operator, which allows for the unification of any same-arity atoms regardless of conflicting symbols.
 [^3] Introducing weak unification greatly increases the search runtime, as one might try to unify any two symbols in the vocabulary at every recursive step.
@@
 The score of the full proof is taken as the minimum of scores across all steps.
-In this work, we apply a similar aggregation; we say that a query fact $s_{1}$ “weakly unifies” with provenance fact $s_{2}$ with unification score equal to the confidence of an NLI model taking $s_{2}$ as the premise and $s_{1}$ as the hypothesis.
-Figure 3: Proposed system framework.
-An off-the-shelf theorem prover searches for proofs of query $\textsc{Prove}(h)$, where symbol $h$ is an NL hypothesis translated from a QA pair.
-The prover uses a set of meta-axioms invoking neural retrieval, entailment, and generation predicates to dynamically instantiate inference rules that use the NL factbase.
+In this work, we apply a similar aggregation; we say that a query fact $s_{1}$ “weakly unifies” with provenance fact $s_{2}$ with a unification score equal to the confidence of an NLI model taking $s_{2}$ as the premise and $s_{1}$ as the hypothesis.
 4 Overview of Approach
-Depicted in Figure 3, our proposed framework is comprised of: An external corpus of facts (some $\{f_{1},\dots f_{n}\}$); a module that converts a QA pair to a hypothesis; an off-the-shelf theorem prover; and a suite of meta-axioms that use neural fact retrieval and dynamic rule generation modules to propose, verify, and score inferences.
-In our experiments, we consider one implementation of this framework that uses the corpus WorldTree (Xie et al., 2020), a set of 9K NL science facts that can interchangeably be considered as rows in 81 $n$-ary relational tables.
-Question Conversion Given a multiple-choice question, the system converts each candidate answer into a hypothesis $h$ using a Question to Declarative Sentence model (QA2D; Demszky et al., 2018) (See §C).
-It then searches for a proof of $h$ against its knowledge base For each alternative, we enumerate $p$ proofs using a time-capped backward chaining search, and then take as the system’s answer the candidate with the overall highest-scoring proof.
+Depicted in Figure 3, our framework is comprised of: an external corpus of facts (some $\{f_{1},\dots f_{n}\}$); a module that converts a QA pair to a hypothesis; an off-the-shelf theorem prover; and a suite of meta-axioms that use neural fact retrieval and dynamic rule generation modules to propose, verify, and score inferences.
+In our experiments, we consider one implementation of this framework that uses the corpus WorldTree Xie et al.
+(2020), a set of 9K NL science facts that can interchangeably be considered as rows in 81 $n$-ary relational tables.
+Question Conversion
+Given a multiple-choice question, the system converts each candidate answer into a hypothesis $h$ using a Question to Declarative Sentence model [QA2D; Demszky et al.
+(2018)] (See §C).
+It then searches for a proof of $h$ against its knowledge base.
+For each alternative, we enumerate $p$ proofs using a time-capped backward chaining search and then take as the system’s answer the candidate with the overall highest-scoring proof.
 4.1 Inference Rule Structure
@@
 Via template I, the system proves the hypothesis by finding a provenance Fact stored in its knowledge store that entails the hypothesis.
-Via template II, it enumerates a pair, Fact1 and Fact2, both either stored in the knowledge store or themselves recursively proved, such that the pair in conjunction entail the hypothesis.
-[^4] Note that a strict two-premise decomposition can prove expressively powerful.
-For example, to prove a hypothesis such as those in EntailmentBank (Dalvi et al., 2021) that require a three-premise conjunction $h\Leftarrow f_{1}\land f_{2}\land f_{3}$, Nellie produces instead a recursive set of decompositions $H\Leftarrow f_{1}\land f_{i};f_{i}\Leftarrow f_{2}\land f_{3}$.
+Via template II, it enumerates a pair, Fact1 and Fact2, both either stored in the knowledge store or themselves recursively proved, such that the pair in conjunction entails the hypothesis.
+[^4] We find that two-premise decomposition is sufficiently powerful and expressive for our purposes.
+For example, to prove a hypothesis such as those in EntailmentBank Dalvi et al.
+(2021) that requires a three-premise conjunction $h\Leftarrow f_{1}\land f_{2}\land f_{3}$, Nellie produces instead a recursive set of decompositions $H\Leftarrow f_{1}\land f_{i};f_{i}\Leftarrow f_{2}\land f_{3}$.
 Template I is given higher search precedence than II, yielding an intuitive high-level procedure: we first look up the hypothesis against the factbase, searching for an entailing fact.
@@
 We define the following core meta-rules, which use the neural predicates Retrieve, Entails, and RuleGen.
-At each step in the backward-chaining search, Nellie’s Prolog engine attempts unify a query with the head of one of these three rules:
+At each step in the backward-chaining search, Nellie’s Prolog engine attempts to unify a query with the head of one of these three rules:
   1. 1.
@@
     Retrieved First Premise Rule Generation
-    $\textsc{Prove}(h)\Leftarrow\textsc{Retrieve}(h^{+},f_{1}^{-})\ \land\ \textsc{RuleGen}(h^{+},f_{1}^{+},f_{2}^{-})$ $\quad\land\textsc{Entails}([f_{1},f_{2}],h)\ \land$ $\textsc{Prove}(f_{2})$
+    $\textsc{Prove}(h)\Leftarrow\textsc{Retrieve}(h^{+},f_{1}^{-})\ \land\ \textsc{RuleGen}(h^{+},f_{1}^{+},f_{2}^{-})$
+    $\quad\land\textsc{Entails}([f_{1},f_{2}],h)\ \land$ $\textsc{Prove}(f_{2})$
 4.2 Unification with Retrieved Facts
@@
 Rule 1 shows how we prove $\textsc{Prove}(h)$ using retrieval.
-The predicate Retrieve proposes candidate $f_{i}$’s given $h$ using a FAISS (Johnson et al., 2019)-based nearest neighbor dense retrieval index.
-[^5] For some situational questions, it is necessary to ground a subquery in evidence from the problem, e.g.
-“A ball is rolling down a hill.” To handle this, we add the question setup (defined as all but its last sentence) as a “fact” always proposed by Retrieve.
+The predicate Retrieve proposes candidate $f_{i}$’s given $h$ using a FAISS Johnson et al.
+(2019)-based nearest neighbor dense retrieval index.
 We train the retrieval encoder via ranking loss such that the embedding for a hypothesis is maximally similar to its supporting facts.
-To promote logical coherence and improve the precision of the system, we apply a set of differently-trained neural RTE models as filters $\textsc{Entails}_{j}(\cdot)$ that iteratively rule out $f_{i}$ candidates that are not classified as entailing $h$.
+To promote logical coherence and improve the precision of the system, we apply a set of neural models for recognizing textual entailment (RTE) as filters $\textsc{Entails}_{j}(\cdot)$ that iteratively rule out $f_{i}$ candidates that are not classified as entailing $h$.
 $\textsc{Entails}(f_{i},h)\Leftarrow\bigwedge_{j=1\dots n}\textsc{Entails}_{j}(f_{i},h)$
 Implicit in rule 1 is that $\textsc{Prove}(h)$ weakly unifies with some $\textsc{Prove}(f_{i})$; we assign the unification score $\theta(h,f_{i})$ equal to the confidence of one RTE model.
+For some questions such as the one depicted in Figure 4, it is necessary to ground a subquery in evidence from the problem.
+To handle this, we add the question setup (defined as all but its last sentence) as a “fact” always proposed by Retrieve.
 4.3 Dynamic Rule Generator (DRG)
-If we do not find an entailing fact for hypothesis $h$, we to decompose $h$ into a pair of entailing premises; we propose candidates using nucleus sampling (Holtzman et al., 2019) from a seq2seq model.
+If we do not find an entailing fact for hypothesis $h$, we decompose $h$ into a pair of entailing premises; we propose candidates using nucleus sampling Holtzman et al.
+(2019) from a seq2seq model.
 The predicate $\textsc{RuleGen}(h,f_{1},f_{2})$ prompts a model trained to generate $h\rightarrow f_{1},f_{2}$ pairs.
-The space of potential decompositions is very large; to bias the proof search towards those more likely to be grounded in the corpus, we adopt a two-pronged approach, illustrated in inference rules 2 and 3, that proposes a heterogeneous search frontier using different biasing strategies.
-Template Conditioned Generation One way in which we might improve our LM-based proposal function is to leverage the high-level structure that supports reasoning in a given domain through template-guided generation, in this case to bias search towards the semi-structure of WorldTree.
+The space of potential decompositions is very large; there are many deductively valid ways to prove a hypothesis in natural language, though only a fraction of them are ultimately groundable in the provided factbase.
+To bias the proof search towards those more likely to be grounded in the corpus, we adopt a two-pronged approach, illustrated in inference rules 2 and 3, that proposes a heterogeneous search frontier using different biasing strategies in addition to straightforward LM sampling.
+Figure 4: Example question in which a sub-hypothesis in the proof tree is grounded to the question context rather than to the factbase.
+Template Conditioned Generation (TCG)
+One way in which we improve our LM-based proposal function is to leverage the high-level structure that supports reasoning in a given domain: we propose template-guided generation to bias search towards the semi-structure of WorldTree.
 [^5] While our approach is applicable to a wide array of reasoning problems, the use of WorldTree-specific templates illustrates one way by which we can infuse domain-specific structure into the neural search algorithm.
-We note that this is a strict departure from the burdensome process of symbolic knowledge engineering; it is instead a simple means by which we can enlist a subject matter expert in order to improve the effectiveness of the engine.
-As our experiments show, this module improves Nellie’s QA accuracy by a few points, though ablating it yields a similarly strong performance.
+This is a strict departure from the burdensome process of symbolic knowledge engineering.
+As our experiments show, this guidance method improves Nellie’s QA accuracy by a few points, though ablating it yields a similarly strong performance.
 WorldTree tables correspond to types of facts with similar syntax and semantics that support scientific reasoning.
-WorldTree QA items are annotated with the facts that support their answers.
+WorldTree questions are annotated with the fact rows that support their answers.
 Each table can be viewed as an n-ary relation of columns whose values are text spans.
-E.g., the Taxonomic relation has columns <A>, [HYPONYM], <is a kind of>, <SCOPE1>, [HYPERNYM], <for>, [PURPOSE].
+E.g., the Taxonomic relation has columns <A>, [HYPONYM], <is a / a kind of>, <SCOPE1>, [HYPERNYM], <for>, [PURPOSE].
 Rows include ‘a bird is an animal’ and ‘a seed is a kind of food for birds.’
-Accordingly, for the RuleGen predicate in rule 2, half of the $f_{1},f_{2}$ candidates are sampled from the DRG conditioned on $h$ plus a template that cues the model to reflect the syntax of a WorldTree table.
+Thus, for the RuleGen predicate in rule 2, half of the $f_{1},f_{2}$ candidates are sampled from the DRG conditioned on $h$ plus a template that cues the model to reflect the syntax of a WorldTree table.
 We train the DRG to accept any masked infilling template (e.g.
-<mask>is a kind of <mask>, akin to those used to pretrain recent LMs (Lewis et al., 2019; Raffel et al., 2020)), and propose decompositions whose first fact reflects the template’s syntax.
+<mask>is a kind of <mask>, akin to those used to pretrain LMs Lewis et al.
+(2019); Raffel et al.
+(2020)), and propose decompositions whose first fact reflects the template’s syntax.
 We thus create a $h,t_{1}\rightarrow f_{1},f_{2}$ model.
 We feed the model templates drawn from WorldTree’s tables, guiding it towards proof steps more likely to be grounded in the factbase.
+A sample of the 150 templates can be found in Figure 5 (a larger list is shown in §D).
 We reuse the same model for non-template-conditioned generation by feeding it an empty template.
-In practice, we make two generation calls: one samples $m$ free-generated candidates, and a second samples $n$ candidates for each of $n_{t}$ templates (see §C).
-Template Selection WorldTree is a diverse corpus, containing tables that are specific to a particular subset of science problems (e.g.
+In practice, we make two generation calls: one samples $m$ free-generated candidates, and a second samples $n$ candidates for each of $n_{t}$ templates.
+WorldTree Relation Template
+KINDOF <mask>is a kind of <mask>
+IFTHEN if <mask>then <mask>
+PROP-THINGS <mask>has <mask>
+CAUSE <mask>causes <mask>
+MADEOF <mask>made of <mask>
+REQUIRES <mask>requires <mask>
+ACTION <mask>is when <mask>
+USEDFOR <mask>used to <mask>
+PARTOF <mask>a part of <mask>
+CHANGE <mask>changes <mask>
+USEDFOR <mask>used for <mask>
+AFFECT <mask>has <mask>impact on <mask>
+PROP-ANIMAL-ATTRIB <mask>is a <mask>animal
+SOURCEOF <mask>is a source of <mask>
+COMPARISON <mask>than <mask>
+EXAMPLES an example of <mask>is <mask>
+COUPLEDRELATIONSHIP as <mask>decreases <mask>will <mask>
+PROP-CONDUCTIVITY <mask>is <mask>conductor
+PROP-GENERIC <mask>is a property of <mask>
+HABITAT <mask>live in <mask>
+MEASUREMENTS <mask>is a measure of <mask>
+Figure 5: Sample of WorldTree templates used for guided generation.
+Template Selection
+WorldTree is a diverse corpus, containing tables that are specific to a particular subset of science problems (e.g.
 the “Predator-Prey” table).
-As conditioning on dozens of templates can be computationally expensive, we introduce a case-based reasoning (Schank, 1983; Das et al., 2021) approach that selects relevant templates for a given hypothesis.
-We construct an Okapi-BM25 (Jones et al., 2000) retrieval index over questions from the WorldTree QA training set to obtain the most lexically similar items to a query.
-At inference time, we select the top-$k$ most similar questions to the query, and take as our template subset the tables of the questions’ annotated facts.
-Retrieval Conditioned Generation In rule 3, rather than generate a pair of subqueries, we immediately ground half of the antecedent by choosing as $f_{1}$ a fact retrieved directly from the corpus.
+As conditioning on dozens of templates can be computationally expensive, we introduce a case-based reasoning Schank (1983); Das et al.
+(2021) approach that selects relevant templates for a given hypothesis.
+We construct an Okapi-BM25 Jones et al.
+(2000) retrieval index over questions from the WorldTree QA training set to obtain the most lexically similar items to a query.
+At inference time, we select the top-$k$ most similar questions to the query and take as our template subset the tables of the questions’ annotated facts.
+Retrieval Conditioned Generation (RCG)
+In rule 3, rather than generate a pair of subqueries, we immediately ground half of the antecedent by choosing as $f_{1}$ a fact retrieved directly from the corpus.
 We have the DRG force decode $f_{1}$ before generating $f_{2}$ as normal, then recur only on $f_{2}$.
-Filters As stochastic sampling from LMs can be a noisy process, some fraction of generated candidates may be invalid: generated facts may be incoherent or not compositionally entail $h$.
-Accordingly, we introduce a suite of compositional entailment verifiers (Khot et al., 2020; Jhamtani and Clark, 2020) trained on two-premise compositional entailment.
-We also add a “self-ask” filter, which following (Tafjord et al., 2022) is an LM fine-tuned to assign a statement a truth value.
-If the confidence is below 0.5 for either an entailment judgement or the ‘self-ask’ belief in $f_{1}$ or $f_{2}$, then we filter the pair.
-All filters condition on the question setup text.
-When Nellie uses these rules, the unification score equals the lowest of the scores for $\textsc{Prove}(f_{1})$, for $\textsc{Prove}(f_{2})$, and the confidence of an entailment filter $s_{e}([f_{1},f_{2}]\Rightarrow H)$.
+Filters
+As stochastic sampling from LMs can be noisy, some fraction of the generated candidate set may be invalid: premises may be incoherent, or the decomposition might not properly entail $h$.
+Accordingly, we introduce a set of compositional entailment verifiers Khot et al.
+(2020); Jhamtani and Clark (2020) trained on two-premise compositional entailment.
+We also add a “self-ask” filter, which following Tafjord et al.
+(2022) is an LM fine-tuned to assign a statement a truth value.
+If the confidence is below 0.5 for either an entailment judgment or the ‘self-ask’ belief in $f_{1}$ or $f_{2}$, then we filter the pair.
+All filters condition on the question text as context.
+When Nellie uses these rules, the unification score equals the lowest of the scores for $\textsc{Prove}(f_{1})$, for $\textsc{Prove}(f_{2})$, and the confidence of entailment filters $s_{e}([f_{1},f_{2}]\Rightarrow H)$.
 4.4 Proof Search
 Given a query, Nellie searches for $t$ seconds to find up to $p$ proofs of depth $d$ or less.
-We follow (Weber et al., 2019) in pruning search branches whose unification score is guaranteed to fall below the current best, given our monotonic aggregation function $\min(\cdot)$.
+We follow Weber et al.
+(2019) in pruning search branches whose unification score is guaranteed to fall below the current best, given our monotonic aggregation function $\min(\cdot)$.
 Found proofs that score under the current best do not count towards $p$.
-Full pseudocode for the algorithm, which follows a depth-first search with a breadth-first lookahead (Stern et al., 2010) to check for unification of generated subgoals, can be found in §E.
+Full pseudocode for the algorithm, which follows a depth-first search with a breadth-first lookahead Stern et al.
+(2010) to check for the unification of generated subgoals, can be found in §E.
+It is parameterized by
+  1. 1.
+    A maximum number of proofs $m$ at which to cut off searching. In experiments, we set this to $10$ for top-level queries and $2$ for recursive subqueries.
+  2. 2.
+    A number of support facts $n_{f}$ to retrieve at each call to $\textsc{Retrieve}_{K}$, which we set to 15.
+  3. 3.
+    Candidate generation rates $n_{v}$ for vanilla nucleus-sampled decompositions, $n_{t}$ for template-conditioned decompositions, and $n_{r}$ for retrieval-conditioned generations.
+We set these each to $40$.
+[^6] Due to batching, these correspond to 3 total calls to the HuggingFace library’s generate function for the T5-3B model.
+Upon removing exact match duplicates, a call to RuleGen produces about $100$ candidates.
+  4. 4.
+    Entailment scoring module $\textsc{Score}_{e}(\cdot)$, which is a separate RTE cross-encoder model for single- and double-premise entailments.
 5 Experiments
@@
 Our experiments illustrate how the approach exemplified by Nellie provides grounded and logical explanations, performing comparably or better than approaches that do not satisfy these properties.
-We evaluate models on two multiple-choice QA datasets constructed so that correct answers are supported by facts in the WorldTree corpus: EntailmentBank (Dalvi et al., 2021) is a dataset of entailment trees for declarativized ARC answers, showing how the hypothesis can be derived via compositional entailment hops from WorldTree facts.
+We evaluate models on two multiple-choice QA datasets constructed so that correct answers are supported by facts in the WorldTree corpus:
+EntailmentBank Dalvi et al.
+(2021) is a dataset of entailment trees for declarativized answers to the AI2 Reasoning Challenge (ARC) dataset Clark et al.
+(2018), showing how the hypothesis can be derived via compositional entailment hops from WorldTree facts.
 We recast the test set, initially designed to test tree reconstruction, into QA by retrieving the corresponding multiple-choice ARC questions from which the hypotheses were constructed.
-WorldTree (Xie et al., 2020) is a subset of the ARC dataset (Clark et al., 2018) annotated with undirected explanation graphs whose nodes are facts from the WorldTree tablestore.
+WorldTree Xie et al.
+(2020) is a subset of the ARC dataset annotated with undirected explanation graphs whose nodes are facts from the WorldTree tablestore.
 We note that WorldTree explanations do not show how facts should combine.
 There is no guarantee that fully grounded trees exist for these questions using the WorldTree corpus alone.
-The difficulty of this task is akin to that of a course exam: the teacher (us) provides the student (the model) a long study guide of facts, but expects the student to use these facts via novel composition in order to reason coherently about a problem.
+The difficulty of this task is akin to that of a course exam: the teacher (us) provides the student (the model) with a very large study guide of facts, but expects the student to use these facts by composing them to reason coherently about a problem.
 Our task metric is accuracy: whether a generated proof of the correct option outscores any other.
@@
 Nellie searches for up to $p$=10 proofs of max depth $d$=5 with a timeout of $t$=180 seconds per option.
-§B discusses further hyperparameters.
-Explanations QA Accuracy (%)
-Grounded Logical Ovr Easy Chal
-EntailmentBank QA
-Nellie (3B) Yes Yes 71.4 76.4 60.4
-Entailer-3B No Yes 48.7 52.8 39.6
- (D $=$ 1) 64.9 71.2 50.9
-Entailer-11B No Yes 73.2 77.3 64.2
- (D $=$ 1) 71.1 76.4 59.4
-WorldTree QA
-Nellie (3B) Yes Yes 71.4 75.7 60.9
-Entailer-3B No Yes 45.2 50.1 33.3
- (D $=$ 1) 47.7 51.6 38.5
-Entailer-11B No Yes 73.2 76.7 64.6
- (D $=$ 1) 74.1 78.7 63.0
-PathNet Yes No 43.4
-TupleILP Yes No 49.8
-ExplanationLP Yes No 62.6
-Diff-Explainer Yes No 71.5
-Table 1: Nellie performance vs. comparable XQA systems.
-Baselines We evaluate Nellie against Entailer (Tafjord et al., 2022), another system that produces entailment tree proofs via backchaining.
+Baselines
+We evaluate Nellie against Entailer Tafjord et al.
+(2022), another system that produces entailment tree proofs via backward chaining.
 Entailer stops recurring not when it finds entailing facts from a corpus, but rather when the model believes that a subquery is true with high confidence.
@@
 To isolate the impact of model size, we also evaluate an Entailer-3B model based on the same T5-3B model as Nellie’s DRG.
-[^8] We obtained the training data from (Tafjord et al., 2022) and trained our model in consultation with the authors.
-We also compare against grounded xQA methods without logical structure (see Fig 2): PathNet (Kundu et al., 2019), which constructs 2-hop chains by linking entities between facts, and three approaches that build graphs from facts using linear programming: TupleILP (Khot et al., 2017), ExplanationLP (Thayaparan et al., 2021), and Diff-Explainer (Thayaparan et al., 2022).
-[^9] We show numbers reported by (Thayaparan et al., 2021, 2022).
-Figure 4: Effect of ablating one or both of rule-conditioned (RCG) and template-conditioned generation (TCG)
-Figure 5: (Left) Example Nellie proofs.
-Top-level queries are decomposed into subqueries via retrieval- or template-conditioned generation.
-Proof leaves are corpus facts.
-(Right) Common classes of error causing false statements to be grounded in true ones.
-Figure 6: Nellie QA accuracy and proof recall on OBQA with vs. without access to common knowledge statements
+[^8] We obtained the training data from Tafjord et al.
+(2022) and trained our model in consultation with the authors.
+We also compare against grounded xQA methods without logical structure (see Fig 2): PathNet Kundu et al.
+(2019), which constructs 2-hop chains by linking entities between facts, and three approaches that build graphs from facts using linear programming: TupleILP Khot et al.
+(2017), ExplanationLP Thayaparan et al.
+(2021), and Diff-Explainer Thayaparan et al.
+(2022).
+[^9] We show results from [Thayaparan et al.
+(2021) & 2022].
+Explanations QA Accuracy (%)
+Grounded Logical Ovr Easy Chal
+EntailmentBank QA
+Nellie (3B) Yes Yes 71.4 76.4 60.4
+Entailer-3B
+(D $=$ 3) No Yes 48.7 52.8 39.6
+(D $=$ 1) No No 64.9 71.2 50.9
+Entailer-11B
+(D $=$ 3) No Yes 73.2 77.3 64.2
+(D $=$ 1) No No 71.1 76.4 59.4
+WorldTree QA
+Nellie (3B) Yes Yes 71.4 75.7 60.9
+Entailer-3B
+(D $=$ 3) No Yes 45.2 50.1 33.3
+(D $=$ 1) No No 47.7 51.6 38.5
+Entailer-11B
+(D $=$ 3) No Yes 73.2 76.7 64.6
+(D $=$ 1) No No 74.1 78.7 63.0
+PathNet Yes No 43.4
+TupleILP Yes No 49.8
+ExplanationLP Yes No 62.6
+Diff-Explainer Yes No 71.5
+Table 1: Nellie performance vs. comparable XQA systems.
 5.1 Results
@@
 Removing both template- and retrieval-conditioned generation lowers Nellie’s performance in all circumstances, highlighting the empirical benefit of structured guidance.
-Ablating TCG (Nellie row vs.
-(-- TCG) and (-- RCG) vs.
-(-- Both)) drops performance by 1-1.5 points.
+Ablating TCG ((-- TCG), and the drop from (-- RCG) to (-- Both)) reduces performance by 1-1.5 points.
 Ablating RCG (Nellie vs (-- RCG)) drops it by 3-4.
@@
 While Nellie was trained on datasets centered around WorldTree, we show that it can perform in another domain by altering the knowledge over which it reasons.
-We consider OpenBookQA (Mihaylov et al., 2018), a dataset that requires reasoning over facts not included in WorldTree; each question is associated with one WorldTree science fact (“metals conduct electricity”), but also one fact from a separate pool of common knowledge (“a suit of armor is made of metal”).
-As with a classical expert system, we have designed Nellie to be “improvable merely by making statements to it” (McCarthy, 1959).
-This means that as we increase its provided knowledge store, we should expect Nellie to reason more effectively.
+We consider OpenBookQA Mihaylov et al.
+(2018), a dataset that requires reasoning over facts not included in WorldTree; each question is associated with one WorldTree science fact (“metals conduct electricity”), but also one fact from a separate pool of common knowledge (“a suit of armor is made of metal”).
+Figure 6: Effect of ablating one or both of rule-conditioned (RCG) and template-conditioned generation (TCG).
+Figure 7: Nellie QA accuracy and proof recall on OBQA with vs. without access to common knowledge statements
+Figure 8: (Left) Example Nellie proofs.
+Top-level queries are decomposed into subqueries via retrieval- or template-conditioned generation.
+Proof leaves are corpus facts.
+(Right) Common classes of error causing false statements to be grounded in true ones.
+Figure 9: Distribution of Nellie error classes among 136 incorrect answers on EntailmentBank questions
+As with a classical expert system, we have designed Nellie to be “improvable merely by making statements to it” McCarthy (1959).
+This suggests that as we increase the provided knowledge store, we should expect Nellie to reason more effectively.
 Because WorldTree does not cover all the knowledge required to answer OBQA questions, we test whether Nellie performs better on them when we add the common knowledge to its retrieval index.
-As the common knowledge annotations are one fact per question and are not designed specifically for entailment, we note that it is unlikely a priori that a fully grounded entailment tree can be created using the available knowledge alone, making this a challenging task for our reasoner.
-Figure 7: Distribution of Nellie error classes among 136 incorrect answers on EntailmentBank questions
+As the common knowledge annotations are one fact per question and are not designed specifically for entailment, it is unlikely a priori that a fully grounded entailment tree can be created using the available knowledge alone, making this a challenging task.
 Figure 7 shows Nellie’s accuracy with and without RCG.
@@
 5.3 Tree Error Analysis
-We find that Nellie can produce high quality proofs of correct hypotheses; examples are shown in Figures 5 (left) and §F.
+We find that Nellie can produce high quality proofs of correct hypotheses; examples are shown in Figure 8 (left) and §F.
 However, the system can also generate proofs for incorrect answers that bypass its entailment filters.
@@
 We list a few categories, depicted in Figure 8 (right).
-A common pattern of error is Hypernym Property Inheritance Errors, in which the model assumes that a member of a taxonomic class has a property of their hypernym, but the property is not universal.
+A common pattern of error is Hypernym Property Inheritance Errors (also known as the “fallacy of the undistributed middle”), in which the model assumes that a member of a taxonomic class has a property of their hypernym, but the property is not universal.
 A colloquial example is inferring penguins can fly from penguins are birds & birds can fly.
-We also observe an amount of Errors from Semantic Drift (Khashabi et al., 2019) between inference hops, culminating in RTE model false positives.
+We also observe an amount of Errors from Semantic Drift Khashabi et al.
+(2019) between inference hops, culminating in RTE model false positives.
 E.g., in block $2$ of Fig 8 (right), which object loses temperature during heat transfer changes between hops.
@@
 The predominant error case is a failure to find any proof of the correct option.
+Correct Answer Tree Analysis
+To investigate whether Nellie reaches a right solution with right vs wrong reasoning, we manually inspected for reasoning errors in 50 correct answer trees produced by NELLIE.
+We found that 39/50 (78%) of trees were perfectly acceptable.
+Most of the 11 unacceptable trees were due to incompleteness at one decomposition step.
 6 Conclusion
-We have proposed a reimagined version of a classical expert system that relies on the inferential power of LLMs rather than handcrafted rules.
+We propose a reimagined version of a classical expert system that relies on the inferential power of LLMs rather than handcrafted rules.
 Nellie has the skeleton of a symbolic theorem prover, but the provided rules invoke neural predicates that allow for systematic reasoning over NL statements.
-In order to dynamically generate inferences, we introduce two mechanisms for knowledge-guided generation: conditioning decompositions on retrieval or model-selected inference templates.
-Our search algorithm undergirds Nellie, the first explainable reasoning system that performs end-to-end QA by searching for grounded entailment trees.
+To dynamically generate inferences, we introduce two mechanisms for knowledge-guided generation: conditioning decompositions on retrieval or model-selected inference templates.
+Our search algorithm undergirds Nellie, an explainable reasoning system that performs end-to-end QA by searching for grounded entailment trees.
 We find that Nellie equals or exceeds the performance of comparable XQA systems, while providing explanations with the simultaneous guarantees of logical directionality and groundedness in human-provided text.
 This work thus suggests a new way to jointly reap the benefits of both modern neural methods and traditional reasoning.
+Acknowledgements
+Thanks to Nils Holzenberger, Kyle Richardson, Elias Stengel-Eskin, Marc Marone, Kate Sanders, Rachel Wicks, and Andrew Blair-Stanek for comments on earlier drafts.
+Thank you to Abbey Coogan for feedback on figure design.
 Appendix A Model Details and Training
@@
 The different neural modules are trained on reformulations of existing datasets for scientific reasoning.
-We use the following sources: WorldTree Explanations [Xie et al., 2020] and EntailmentBank [Dalvi et al., 2021] are used as described in the paper.
-eQASC [Jhamtani and Clark, 2020] and SciTaiL [Khot et al., 2018] are datasets of two- and one-premise entailment questions.
-QA2D Our QA2D model is a T5-11B [Raffel et al., 2020] fine-tuned by [Tafjord et al., 2022] on the original QA2D dataset [Demszky et al., 2018].
-Retrieval Module The dense retrieval model is a SentenceTransformers Siamese BERT-Network encoder [Reimers and Gurevych, 2019] pretrained on the MS MARCO IR corpus [Nguyen et al., 2016], which we fine-tune via ranking loss to maximize the cosine similarity between a hypothesis and its support facts.
+We use the following sources: WorldTree Explanations Xie et al.
+[2020] and EntailmentBank Dalvi et al.
+[2021] are used as described in the paper.
+eQASC Jhamtani and Clark [2020] and SciTaiL Khot et al.
+[2018] are datasets of two- and one-premise entailment questions.
+QA2D Our QA2D model is a T5-11B Raffel et al. [2020] fine-tuned by Tafjord et al. [2022] on the original QA2D dataset Demszky et al. [2018].
+Retrieval Module
+The dense retrieval model is a SentenceTransformers Siamese BERT-Network encoder Reimers and Gurevych [2019] pretrained on the MS MARCO IR corpus Nguyen et al.
+[2016], which we fine-tune via ranking loss to maximize the cosine similarity between a hypothesis and its support facts.
 We gather $h,f_{1}$ from the WorldTree, EntailmentBank, and eQASC training sets, then sample random negatives during training.
-Dynamic Rule Generator The DRG is a T5-3B model [Raffel et al., 2020] fine-tuned on $h\rightarrow[f_{1},f_{2}]$ pairs drawn from eQASC, which are numerous (26K) but noisy, and the binary composition steps from the EntailmentBank training set, which are few (3.8K) but ultimately rooted in facts from WorldTree.
-To make the DRG able to condition on templates, we adopt the span masking strategy from [Raffel et al., 2020] to create a randomly masked version $\hat{f}_{1}$ of each $f_{1}$ and then train the model on $h,\hat{f}_{1}\rightarrow f_{1},f_{2}$ pairs.
+Dynamic Rule Generator
+The DRG is a T5-3B model Raffel et al.
+[2020] fine-tuned on $h\rightarrow[f_{1},f_{2}]$ pairs drawn from eQASC, which are numerous (26K) but noisy, and the binary composition steps from the EntailmentBank training set, which are few (3.8K) but ultimately rooted in facts from WorldTree.
+To make the DRG able to condition on templates, we adopt the span masking strategy from Raffel et al.
+[2020] to create a randomly masked version $\hat{f}_{1}$ of each $f_{1}$ and then train the model on $h,\hat{f}_{1}\rightarrow f_{1},f_{2}$ pairs.
 For each original triplet, we create one training example with a template, and one in which $\hat{f}_{1}$ is an empty <mask>, giving the model the dual capacity for non-template-generation.
-Entailment Filters We find that a mixture of strategies helps prioritize precision over recall when filtering candidates.
+Entailment Filters
+We find that a mixture of strategies helps prioritize precision over recall when filtering candidates.
 For 1-premise filtering for unification and 2-premise filtering for rule generation, we fine-tune a separate pair of models consisting of a SentenceTransformer cross-encoder classifier and a T5 seq2seq.
@@
 The 1-premise filters are trained on the SciTail dataset, the 2-premise filters on eQASC and EntailmentBank.
-For 2-premise entailment, we additionally add a “self-ask” filter, which following [Tafjord et al., 2022] is a T5 model fine-tuned to produce its 0-to-1 confidence that a statement is true.
+For 2-premise entailment, we additionally add a “self-ask” filter, which following Tafjord et al.
+[2022] is a T5 model fine-tuned to produce its 0-to-1 confidence that a statement is true.
 If the confidence is below 0.5 for either $f_{1}$ or $f_{2}$, we filter the pair.
-As in [Tafjord et al., 2022], we train our DRG seq2seq in a multitask fashion so that it serves as both the rule generator as well as the “self-ask” filter and (half of) the 2-premise entailment filter.
+As in Tafjord et al.
+[2022], we train our DRG seq2seq in a multitask fashion so that it serves as both the rule generator as well as the “self-ask” filter and (half of) the 2-premise entailment filter.
 A.1 Training
@@
 […]
-We used Optuna [Akiba et al., 2019] to find the number of facts for retrieval-conditioned generation, the $p$ value for nucleus sampling, the entailment threshold, and the number of candidates generated at each recursion step.
+We used Optuna Akiba et al.
+[2019] to find the number of facts for retrieval-conditioned generation, the $p$ value for nucleus sampling, the entailment threshold, and the number of candidates generated at each recursion step.
 We used as our objective the overall QA accuracy on the EntailmentBank dev set.
@@
 Appendix E Search Algorithm
-Algorithm 1 depicts Nellie’s search algorithm, the logic of which is implemented in DeepProbLog [Manhaeve et al., 2018]. The search is parametrized by
-  1. 1.
-    A maximum number of proofs $m$ at which to cut off searching. We set this to $10$ for top-level queries and $2$ for recursive subqueries.
-  2. 2.
-    Candidate generation rates $n_{v}$ for vanilla nucleus-sampled decompositions, $n_{t}$ for template-conditioned decompositions, and $n_{r}$ for retrieval-conditioned generations.
-We set these to $200$, $250$ and $200$, respectively.
-[^11] Due to batching, these correspond to 1, 2, and 1 total calls to HuggingFace’s generate function.
-Upon removing exact match duplicates, a call to RuleGen produces about $400$ candidates.
-  3. 3.
-    The number of support facts $n_{f}$ to retrieve at each call to $\textsc{Retrieve}_{K}$, which we set to 15.
-  4. 4.
-    Entailment scoring module $s_{e}(\cdot)$, which is a separate NLI cross-encoder model for single- and double-premise entailments.
-1:procedure Prove($H$, $m$) $\rightarrow$ proofs $tree^{(1)}(H),\dots tree^{(m)}(H)$ scores $s^{(1)}(H),\dots s^{(m)}{H}$:
+Algorithm 1 depicts Nellie’s search algorithm, the logic of which is implemented in a meta-interpreter coded in DeepProbLog Manhaeve et al. [2018].
+Appendix F Example Proofs
+Figure 11 illustrates more proofs generated by Nellie for questions from EntailmentBank and WorldTree.
+Appendix G Impact of Design Choices
+Nellie is comprised of numerous modules that are very difficult to evaluate in isolation given a lack of in-domain test sets for tasks like QA2D, decomposition generation ($H\rightarrow F1,F2$) compositional RTE ($H\Leftarrow F1,F2$), single-premise RTE ($H\Leftarrow F1$), and fact retrieval.
+Below, we enumerate some modeling choices, alternatives we tried, and the extent to which we observed them impact the system.
+  • QA2D:
+    Nellie uses:
+    Tafjord et al. [2022]’s T5-11B model.
+    We also tried: using GPT-J Wang and Komatsuzaki [2021] with in-context-learning exemplars drawn from EntailmentBank.
+This results in 10+ point drop in QA, mainly because (A) EntailmentBank hypotheses are designed to be generics instead of question-conditioned statements and (B) GPT-J tended to generate true statements even for wrong answer options.
+  • Decomposition generation ($H\rightarrow F1,F2$):
+    Nellie uses: A T5-3B model trained in a multi-angle fashion similar to Tafjord et al. [2022], using additional data for compositional entailment and decomposition generation.
+    We also tried: Using a T5-Large model trained only on the decomposition data.
+This allowed us to generate far more decomposition candidates per hypothesis, but rate of obtaining a good decomposition was far lower.
+This led to 5+ point drop in QA.
+  • Compositional RTE ($H\Leftarrow F1,F2$):
+    Nellie uses:
+    A sequence of (1) The same T5-3B model as for generation and (2) a cross-encoder trained on eQASC and EntailmentBank.
+    We also tried: (A) using a bi-encoder trained via contrastive loss (this was not as effective at ruling out bad decompositions), (B) not training the bi-encoder to condition on the question, and (C) only training on EntailmentBank (as Tafjord et al.
+[2022] do).
+All three variants reduce QA by 3-7%.
+  • Single-premise RTE ($H\Leftarrow F1$):
+    Nellie uses:
+    (1) A cross-encoder trained on Sci-Tail
+    We also tried: an off-the-shelf RTE model [^11] https://huggingface.co/cross-encoder/nli-deberta-v3-base, which had too high a false positive rate to be effective.
+  • Fact Retrieval:
+    Nellie uses:
+    a Sentence-Transformers bi-encoder pretrained on MS MARCO and fine-tuned on $(H,F1)$ pairs drawn from WorldTree, EntailmentBank, and eQASC.
+    We also tried: the same bi-encoder off-the-shelf without finetuning. This only dropped performance by 3-4 points.
+1:procedure Prove($H$, $m$, $d$) $\rightarrow$ proofs $tree^{(1)}(H),\dots tree^{(m)}(H)$ scores $s^{(1)}(H),\dots s^{(m^{\prime})}{(H)}$:
 2:   i. check weak unification against $K$:
@@
 […]
-6:   ii. generate candidate decompositions: // If $H$ does not unify against $K$, decompose it
-7:       decompositions $D={\color[rgb]{0,0,1}\textsc{RuleGen}}(H)$
-8:   iii. check whether any decompositions pairs both unify: // this step amortizes premise lookups
-9:       proofs = []; scores = []
-10:       forall $(p_{1},p_{2})\in D$ do
-11:               if $p_{1}\in K$  and  $\exists p_{2}^{\prime}\in{\color[rgb]{0,0,1}\textsc{PremiseLookup}}(p_{2})$ // if $p_{1}$ already grounded, its subtree is leaf
-12:               then proofs += ($(\{p_{1},(p_{2}^{\prime}\vdash{}p_{2})\}\vdash{}H)$
-13:                   scores += $\min({\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{2}^{\prime}\vdash{}p_{2}),{\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{1}\land p_{2}\vdash{}H))$ // tree score is min of scores used to construct it
-14:               elif $\exists p_{1}^{\prime}\in{\color[rgb]{0,0,1}\textsc{PremiseLookup}}(p_{1})$  and  $\exists p_{2}^{\prime}\in{\color[rgb]{0,0,1}\textsc{PremiseLookup}}(p_{2})$
-15:               then proofs += ($(\{(p_{1}^{\prime}\vdash{}p_{1}),(p_{2}^{\prime}\vdash{}p_{2})\}\vdash{}H)$
-16:                   scores += $\min({\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{1}^{\prime}\vdash{}p_{1}),{\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{2}^{\prime}\vdash{}p_{2}),{\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{1}\land p_{2}\vdash{}H))$
-17:       if proofs not empty
-18:       then return proofs, scores
-19:   iv. recur on generated premises: // if still no proof, need to recur
-20:       proofs = []; scores = []
-21:       forall $(p_{1},p_{2})\in D$ do
-22:               if ${\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{1}\land p_{2}\vdash{}H)>\max(\text{scores})$  and  $\text{len}(\text{proofs})<m$ // only recur if chance of better score
-23:               then $(\text{proof}_{p_{1}},S_{p_{1}})={\color[rgb]{0,0,1}\textsc{Prove}}(p_{1},1)$
-24:                   $(\text{proof}_{p_{2}},S_{p_{2}})={\color[rgb]{0,0,1}\textsc{Prove}}(p_{2},1)$
-25:                   recursive score $S_{r}=\min(S_{p_{1}},S_{p_{2}},{\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{1}\land p_{2}\vdash{}H))$
-26:                   if $S_{r}>\max(\text{scores})$
-27:                   then proofs += ($(\{\text{proof}_{p_{1}},\text{proof}_{p_{2}}\}\vdash{}H)$
-28:                      scores += $S_{r}$
-29:       return proofs, scores
-30:
-31:procedure PremiseLookup(H) $\rightarrow$ premises $P$ s.t. $\forall p\in P,p\vdash{}H$ // try to ground H in K via entailment
-32:       candidate premises $P={\color[rgb]{1,.5,0}\textsc{Retrieve}_{K}}(H,n_{f})$ // retrieve candidate premises s.t. $p\vdash{}H$
+6:       elif $d=d_{\text{max}}$
+7:       then Fail
+8:   ii. generate candidate decompositions: // If $H$ does not unify against $K$, decompose it
+9:       decompositions $D={\color[rgb]{0,0,1}\textsc{RuleGen}}(H)$
+10:   iii. check whether any decompositions pairs both unify: // this step amortizes premise lookups
+11:       proofs = []; scores = []
+12:       forall $(p_{1},p_{2})\in D$ do
+13:               if $p_{1}\in K$  and  $\exists p_{2}^{\prime}\in{\color[rgb]{0,0,1}\textsc{PremiseLookup}}(p_{2})$ // if $p_{1}$ already grounded, its subtree is leaf
+14:               then proofs += ($(\{p_{1},(p_{2}^{\prime}\vdash{}p_{2})\}\vdash{}H)$
+15:                   scores += $\min({\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{2}^{\prime}\vdash{}p_{2}),{\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{1}\land p_{2}\vdash{}H))$ // tree score is min of scores used to construct it
+16:               elif $\exists p_{1}^{\prime}\in{\color[rgb]{0,0,1}\textsc{PremiseLookup}}(p_{1})$  and  $\exists p_{2}^{\prime}\in{\color[rgb]{0,0,1}\textsc{PremiseLookup}}(p_{2})$
+17:               then proofs += ($(\{(p_{1}^{\prime}\vdash{}p_{1}),(p_{2}^{\prime}\vdash{}p_{2})\}\vdash{}H)$
+18:                   scores += $\min({\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{1}^{\prime}\vdash{}p_{1}),{\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{2}^{\prime}\vdash{}p_{2}),{\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{1}\land p_{2}\vdash{}H))$
+19:       if proofs not empty
+20:       then return proofs, scores
+21:   iv. recur on generated premises: // if still no proof, need to recur
+22:       proofs = []; scores = []
+23:       forall $(p_{1},p_{2})\in D$ do
+24:               if ${\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{1}\land p_{2}\vdash{}H)>\max(\text{scores})$  and  $\text{len}(\text{proofs})<m$ // only recur if chance of better score
+25:               then $(\text{proof}_{p_{1}},S_{p_{1}})={\color[rgb]{0,0,1}\textsc{Prove}}(p_{1},1,d+1)$
+26:                   $(\text{proof}_{p_{2}},S_{p_{2}})={\color[rgb]{0,0,1}\textsc{Prove}}(p_{2},1,d+1)$
+27:                   recursive score $S_{r}=\min(S_{p_{1}},S_{p_{2}},{\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{1}\land p_{2}\vdash{}H))$
+28:                   if $S_{r}>\max(\text{scores})$
+29:                   then proofs += ($(\{\text{proof}_{p_{1}},\text{proof}_{p_{2}}\}\vdash{}H)$
+30:                      scores += $S_{r}$
+31:       return proofs, scores
+32:
+33:procedure PremiseLookup(H) $\rightarrow$ premises $P$ s.t. $\forall p\in P,p\vdash{}H$ // try to ground H in K via weak entailment
+34:       candidate premises $P={\color[rgb]{1,.5,0}\textsc{Retrieve}_{K}}(H,n_{f})$ // retrieve candidate premises s.t. $p\vdash{}H$
 […]
-34:               $P=f(P,H)$ // only keep premises that pass all filters
-35:       return $P$
-36:
-37:procedure RuleGen(H) $\rightarrow$ decompositions $D=(p_{1}^{1},p_{2}^{1}),\dots(p_{1}^{c},p_{2}^{c})$ // find candidate pairs s.t. $p_{1}\land p_{2}\vdash{}H$
-38:   D = []
-39:   $D$ += ${\color[rgb]{1,.5,0}\textsc{Generate}}(H,n=n_{v})$
-40:   WTTemplates = CaseBasedRetrieveTemplate(H)
-41:   outputs per template $o_{t}=n_{t}/\text{len}({\color[rgb]{0.56,1,0.26}\textsc{WTTemplates}})$
-42:   forall $t\in{\color[rgb]{0.56,1,0.26}\textsc{WTTemplates}}$ do // generate candidates for each of 50 templates in WTTemplates
-43:       $D$ += ${\color[rgb]{1,.5,0}\textsc{Generate}}([H\ t],n=o_{t})$ // T5 model accepts optional $t$ appended to $H$
-44:   retrieved first premises $P_{1}={\color[rgb]{1,.5,0}\textsc{Retrieve}_{K}}(H,n_{f})$
-45:   outputs per premise $o_{r}=n_{r}/\text{len}(P_{1})$
-46:   forall $p_{K}\in P_{1}$ do // generate candidates with retrieved support facts as first premise
-47:         $D$ += ${\color[rgb]{1,.5,0}\textsc{Generate}}(H,\text{prefix}=p_{K},n=o_{r})$ // Force decode $p_{K}$ then generate a $p_{2}$
+36:               $P=f(P,H)$ // only keep premises that pass all filters
+37:       return $P$
+38:
+39:procedure RuleGen($H$) $\rightarrow$ decompositions $D=(p_{1}^{1},p_{2}^{1}),\dots(p_{1}^{c},p_{2}^{c})$ // find candidate pairs s.t. $p_{1}\land p_{2}\vdash{}H$
+40:       $D$ = []
+41:       $D$ += ${\color[rgb]{1,.5,0}\textsc{Generate}}(H,n=n_{v})$
+42:       WTTemplates = CaseBasedRetrieveTemplate(H)
+43:       outputs per template $o_{t}=n_{t}/\text{len}({\color[rgb]{0.51953125,0.69921875,0.421875}\textsc{WTTemplates}})$
+44:       forall $t\in{\color[rgb]{0.51953125,0.69921875,0.421875}\textsc{WTTemplates}}$ do // generate candidates for each of 50 templates in WTTemplates
+45:               $D$ += ${\color[rgb]{1,.5,0}\textsc{Generate}}([H\ t],n=o_{t})$ // T5 model accepts optional $t$ appended to $H$
+46:       retrieved first premises $P_{1}={\color[rgb]{1,.5,0}\textsc{Retrieve}_{K}}(H,n_{f})$
+47:       outputs per premise $o_{r}=n_{r}/\text{len}(P_{1})$
+48:       forall $p_{K}\in P_{1}$ do // generate candidates with retrieved support facts as first premise
+49:               $D$ += ${\color[rgb]{1,.5,0}\textsc{Generate}}(H,\text{prefix}=p_{K},n=o_{r})$ // Force decode $p_{K}$ then generate a $p_{2}$
 […]
-49:               $D=f(D,H)$ // only keep premise pairs that pass all filters
-50:         return $D$
-Algorithm 1 Nellie’s backchaining algorithm, coded in Prolog, for searching for proof list $tree^{(1)}(H),\dots tree^{(m)}(H)$ with scores $s^{(1)}(H),\dots s^{(m)}(H)$ for a hypothesis $H$ against factbase $K$.
-The rule generator produces $n_{v}$ candidates via vanilla nucleus sampling, $n_{t}$ candidates via sampling from templates, and $n_{r}$ candidates whose first fact is retrieved from K.
-All Generate, Retrieve, and entailment scorer $\textsc{Score}_{e}$ calls in forall loops are batched.
-All module outputs are cached.
-OnePremiseFilters and TwoPremiseFilters  are ordered lists of neural entailment models.
-Appendix F Example Proofs
-Figure 9 illustrates more proofs generated by Nellie for questions from EntailmentBank and WorldTree.
-Blue-outlined nodes are hypotheses generated by the QA2D model for the correct answer highted in green.
-Grey nodes are generated subqueries; tan nodes are leaves from Nellie’s knowledge corpus.
-The green node is a question setup text that grounds a contextual subquery via entailment.
-Question
-Nellie Proof
-Cheetahs have come close to extinction due to hunting, drought, and disease.
-There is now very little genetic variation in cheetah populations.
-Which of the following is a result of the limited genetic variation in the current cheetah populations compared to earlier cheetah populations with more variation?
-(A) Cheetahs in current populations are more resistant to new diseases.
-(B) The survival rate of young cheetahs is increased in current populations.
-(C) Cheetahs in current populations are less able to interbreed with other species.
-(D) The current cheetah populations are less likely to be able to adapt to environmental changes.
-In a city near the ocean, fog often forms on summer mornings.
-Which of the following statements best explains how this fog forms?
-(A) Ocean water evaporates and then condenses in the air.
-(B) Crashing waves spray tiny drops of ocean water into the air.
-(C) Water runoff moves toward the ocean and collects near the shore.
-(D) Rain clouds move in from the ocean and evaporate as they reach the shore.
+51:               $D=f(D,H)$ // only keep premise pairs that pass all filters
+52:       return $D$
+[…]
+Question Nellie Proof
+In what way can forest fires affect the lithosphere? (A) increasing soil erosion (B) causing air pollution (C) increasing seismic activity (D) causing violent dust storms Refer to caption
 The snowshoe hare sheds its fur twice a year.
@@
 (D) Shedding fur helps the hare blend into its habitat.
+Refer to caption
 Scientists use large optical telescopes to obtain information about the planets in the solar system.
 What wavelengths of electromagnetic radiation provide this information?
-(A) gamma radiation (B) infrared radiation (C) radio waves (D) visible light
+(A) gamma radiation (B) infrared radiation (C) radio waves (D) visible light Refer to caption
 Figure 11: Example multiple-choice questions with NELLIE’s answer and corresponding proof.
+Blue-outlined nodes are hypotheses generated by the QA2D model for the correct answer highted in green.
+Grey nodes are generated subqueries; tan nodes are leaves from Nellie’s knowledge corpus.